### v3

# A Framework of SO(3)-equivariant Non-linear Representation Learning and its Application to Electronic-Structure Hamiltonian Prediction

Abstract
We present both a theoretical and a methodological framework that addresses a critical challenge in applying deep learning to physical systems: the reconciliation of non-linear expressiveness with SO(3)-equivariance in predictions of SO(3)-equivariant quantities. Inspired by covariant theory in physics, we address this problem by exploring the mathematical relationships between SO(3)-invariant and SO(3)-equivariant quantities and their representations. We first construct theoretical SO(3)-invariant quantities derived from the SO(3)-equivariant regression targets, and use these invariant quantities as supervisory labels to guide the learning of high-quality SO(3)-invariant features. Given that SO(3)-invariance is preserved under non-linear operations, the encoding process for invariant features can extensively utilize non-linear mappings, thereby fully capturing the non-linear patterns inherent in physical systems. Building on this foundation, we propose a gradient-based mechanism to induce SO(3)-equivariant encodings of various degrees from the learned SO(3)-invariant features. This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations, while theoretically preserving their equivariant properties as we prove. We apply our theory and method to the electronic-structure Hamiltonian prediction tasks, experimental results on eight benchmark databases covering multiple types of elements and challenging scenarios show dramatic breakthroughs on the state-of-the-art prediction accuracy, with improvements of up to 40% in predicting Hamiltonians and up to 76% in predicting downstream physical quantities such as occupied orbital energy. Our approach goes beyond handling physical systems and offers a promising general solution to the critical dilemma between equivariance and non-linear expressiveness for the deep learning paradigm.
keywords:
Deep Learning, Neural Networks, SO(3)-equivariance, Atomic Modeling, Hamiltonian Prediction
1 Introduction
With the advantages in computational complexity and generalization capabilities, deep learning paradigm has vigorously driven advancements in physics research [1]. For example, in predicting electronic-structure Hamiltonians, traditional Density Functional Theory (DFT) methods [2, 3] suffer from computational complexities of $\mathcal{O}(N^{3})$, where $N$ is the number of atoms in a system. In contrast, deep learning approaches have significantly reduced this complexity to $\mathcal{O}(N)$ [4, 5, 6, 7, 8], accelerating processes by thousands or even millions of times. This reduction in computational demand not only unlocks possibilities for analyzing extremely large atomic systems but also paves the way for efficient material simulation, discovery, and design that were previously unimaginable.
However, deep learning methods still face substantial challenges when processing physical systems. To align with fundamental physical laws, these methods must strictly adhere to symmetry principles. For example, physical quantities such as force fields and electronic-structure Hamiltonians must be equivariant under 3D rotational transformations, i.e., elements from the SO(3) group. Traditional deep learning methods [9, 10] typically approximated SO(3)-equivariance through rotational data augmentation. However, this approach does not provide the strict assurance of SO(3)-equivariance required for physical research. Besides, the calculation of physical quantities calls for high numerical accuracy, necessitating that the neural networks possess strong capabilities to express the complex non-linear mappings from atomic structures to the regression targets. However, existing deep learning methods are unable to simultaneously ensure strict SO(3)-equivariance and high numerical accuracy. The root cause of this problem lies in the conflict between SO(3)-equivariance and non-linear expressiveness: specifically, directly applying non-linear activation functions on SO(3)-equivariant features (with degree $l\geq 1$) may lead to the loss of equivariance, while bypassing non-linear mappings severely restricts the network’s expressive capabilities and thereby lowering down the achievable accuracy. As pointed out in previous literature [11], this issue is commonly found in machine learning tasks that demand both strict equivariance and fine-grained generalization performance. Although several efforts have been made to alleviate this issue by introducing gated activation mechanisms [8, 7, 12, 13, 14], developing local coordinate strategy [6], and employing discrete spherical channels [11, 15, 16], they still, to varying degrees, compromised either non-linear expressiveness [7, 8, 12, 13, 14] or strict SO(3)-equivariance [11, 15, 16] when processing physical systems. Please refer to Supplementary Information A for an in-depth analysis.
To address the equivariance-expressiveness dilemma, we make theoretical and methodological explorations on unifying strict SO(3)-equivariance with strong non-linear expressiveness within the realm of deep representation learning. We are inspired by the insight that invariant quantities in transformation often reflect the mathematical nature of physical laws and can induce other quantities with equivariant properties. For example, in special relativity [18], the spacetime interval between events is an invariant under Lorentz transformations. This invariance is foundational in formulating physical laws that involve covariant quantities. Furthermore, from the perspective of deep representation learning, the attribute that invariance is preserved under non-linear operations is a significant advantage, given its compatibility with non-linear expressive capabilities. Built upon these insights, we theoretically solve the equivariance-expressiveness dilemma by intensively exploring and making use of the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities and representations: we first dedicate efforts to learning high-quality SO(3)-invariant features with ample non-linear expressiveness, and subsequently, we derive SO(3)-equivariant non-linear representations from these SO(3)-invariant ones. Specifically:
First, we construct quantities that are theoretically SO(3)-invariant, namely $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$, where $\mathbf{Q}$ are the SO(3)-equivariant regression targets, $tr(\cdot)$ signifies the trace operation, and $\dagger$ denotes the conjugate transpose operation. A significant advantage of these SO(3)-invariant quantities lies in the fact that they are directly derived from the SO(3)-equivariant target labels and can serve as unique supervision labels for the effective learning of informative SO(3)-invariant features that capture the intrinsic symmetry properties of the mathematical structure of $\mathbf{Q}$ without requiring additional labeling resources.
Second, we propose a gradient-based mechanism to induce SO(3)-equivariant representations for predicting the regression target $\mathbf{Q}$, from high-quality non-linear SO(3)-invariant features learned under the supervision of $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$. This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations while preserving their equivariant properties, as we prove.
We develop our theory into a unified SO(3)-equivariant non-linear representation learning method, and apply it to the computation of electronic-structure Hamiltonian, which occupies a central position in quantum mechanics to induce key physical quantities such as electronic wavefuction, orbital energy, and band structure, bringing deeply understanding of the electrical, optical, magnetic, and transport properties of atomic systems, while posing significant difficulties for machine learning techniques due to its intrinsic symmetry with respect to the SO(3) group as well as the high-dimensional complexity, as shown in Fig. 1. Our method dramatically enhances the state-of-the-art performance on eight benchmark Hamiltonian databases from the renowned DeepH and QH9 benchmark series [6, 8, 19], showing excellent generality to various challenging scenarios such as thermal motions, bilayer twists, scale variations, and new trajectories, with an accuracy improvement of Hamiltonians of up to 40%. Moreover, our approach substantially enhances the prediction accuracy of downstream physical quantities of Hamiltonians including occupied orbital energies and electronic wavefunctions, with an accuracy improvement by a maximum of 76% and 17%, respectively. Our leading performance comprehensively demonstrate that our method, while satisfying SO(3)-equivariance, possesses excellent expressive power and generalization performance, providing an effective deep learning tool for efficient and accurate electronic-structure calculations of atomic systems.
Figure 1: Illustration of the high-dimensional complexity of the Hamiltonian. In an atomic system, a Hamiltonian matrix is defined between each pair of atoms within each other’s cutoff radius, constituting a large cluster of Hamiltonian matrices for all atom pairs. Besides, each Hamiltonian matrix contains multiple basic blocks, with the $\mathbf{Q}=\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 2 corresponding to a basic block $\mathbf{H}^{l_{p}\otimes l_{q}}$ here. Thus, for a Hamiltonian prediction model, the space of the regression target is typically huge.
2 Problem Formalization
Physical quantities that are SO(3)-equivariant typically possess their respective equivariant formulas under both matrix-formed direct-product states and vector-formed direct-sum states, which are mathematically equivalent to each other.
Without loss of generality, let ${\bf{Q}}^{l_{p}\otimes l_{q}}$ denote an SO(3)-equivariant quantity in direct-product state formed by $l_{p}\otimes l_{q}$, i.e., the direct product between degrees $l_{p}$ and $l_{q}$. It obeys the following SO(3)-equivariant law:
[…]
where $\dagger$ denotes the conjugate transpose operation. ${\bf{R}}\in\mathbb{R}^{3\times 3}$ is the rotational matrix and ${\bf{D}}^{l_{p}}({\bf{R}})\in\mathbb{R}^{(2l_{p}+1)\times(2l_{p}+1)}$, ${\bf{D}}^{l_{q}}({\bf{R}})\in\mathbb{R}^{(2l_{q}+1)\times(2l_{q}+1)}$ are the Wigner-D matrices of degrees $l_{p}$ and $l_{q}$, respectively, ${\bf{Q}}({\bf{R}})^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ denotes the transformed results of ${\bf{Q}}^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ through the rotational operation by ${\bf{R}}$.
${\bf{Q}}^{l_{p}\otimes l_{q}}$ in the direct-product state can be further decomposed into a series of direct-sum state components with vector forms, i.e., ${\bf{q}}^{l}(|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q})$, which follows SO(3)-equivariant law mathematically equivalent to Eq. (1) but with a simpler form:
${\bf{q}}({\bf{R}})^{l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{q}}^{l},|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q}$ (2)
where ${\bf{q}}^{l}\in\mathbb{R}^{2l+1}$ and ${\bf{q}}({\bf{R}})^{l}\in\mathbb{R}^{2l+1}$ respective denote the components with degree $l$ before and after the rotational operation by ${\bf{R}}$.
Eq. (1) and Eq. (2) are mathematical equivalent to each other, meaning that fulfilling one under the corresponding state naturally satisfies the other. Due to the simplicity and ease of handling of the direct-sum state form, existing deep learning methods [6, 8, 7] typically first predict $\mathbf{q}^{l}(|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q})$, then convert it to $\mathbf{Q}^{l_{p}\otimes l_{q}}$ through Clebsch-Gordan matrices. In the direct-sum state, to obey the equivariant law of Eq. (2) for the regression target, the hidden representations of neural networks must also satisfy the same form of equivariance:
${\bf{f}}({\bf{R}})^{(k)l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{f}}^{(k)l}$ (3)
where ${\bf{f}}^{(k)l}\in\mathbb{R}^{2l+1}$ and ${\bf{f}}({\bf{R}})^{(k)l}\in\mathbb{R}^{2l+1}$ respectively denote one channel of features with degree $l$ before and after the rotational operation by ${\bf{R}}$, at the $k$ th hidden layer.
Due to the intrinsic complexity and non-linearity of physic quantities, neural networks on regressing these quantities are supposed to equip with non-linear mappings denoted as $g_{nonlin}(\cdot)$ to fully capture the intrinsic patterns of the physical quantities, which is crucial for precise and generalizable prediction performance. Meanwhile, $g_{nonlin}(\cdot)$ must also preserve SO(3)-equivariance, which is expressed as:
${\bf{f}}({\bf{R}})^{(k+1)l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{f}}^{(k+1)l},{\bf{f}}^{(k+1)l}=g_{nonlin}({\bf{f}}^{(k)l})$ (4)
However, directly implementing $g_{nonlin}(\cdot)$ as neural network module with non-linear activation functions, such as $Sigmoid$, $Softmax$ and $Silu$, may result in the destruction of strict equivariance. How to make $g_{nonlin}(\cdot)$ both theoretically SO(3)-equivariant and capable of non-linear expressiveness, is the core problem this paper aims at solving.
3 Theory
This section establishes a rigorous theoretical foundation for integrating SO(3)-equivariance with non-linear expressiveness in the deep learning paradigm. First, we propose a theoretical construct to derive SO(3)-invariant quantities from the SO(3)-equivariant targets, aiming at facilitating the learning of high-quality SO(3)-invariant features. Second, we propose a gradient-based mechanism encoding features maintaining theoretically SO(3)-equivariance while possessing the non-linear expressive capability induced from SO(3)-invariant non-linear features to capture the complex non-linear patterns inherent in physical quantities.
Theorem 1.
The quantity $\mathbf{T}=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is SO(3)-invariant, where $\mathbf{Q}$ is the simplified representation (without superscripts) of $\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 2, and $\dagger$ denotes the conjugate transpose operation, $tr(\cdot)$ is the trace operation.
Theorem 2.
The non-linear neural mapping defined as the following is SO(3)-equivariant, satisfying the requirements specified in Eq. (4):
$\mathbf{v}=g_{nonlin}(\mathbf{f})=\frac{\partial z}{\partial\mathbf{f}}\quad\text{s.t.}\quad z=s_{nonlin}(u),u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ (5)
where $\mathbf{f}$ is an input SO(3)-equivariant feature with degree $l$ in the direct-sum state, $\otimes$ denotes the direct-product operation of tensors, $\text{CGDecomp}(\cdot,0)$ refers to performing a Clebsch-Gordan decomposition of the tensor and returning the scalar component of degree 0, $s_{nonlin}(\cdot)$ represents non-linear neural modules, and $\mathbf{v}$ is the outputted feature encoded by $g_{nonlin}(\cdot)$.
Proofs of the two theorems are included in Supplementary Information B.
Remark 1.
In Theorem 1, the term $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is equal to the square of the Frobenius norm of the matrix $\mathbf{Q}$, i.e., $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\|\mathbf{Q}\|_{F}^{2}=\sum_{m,n}|\mathbf{Q}_{mn}|^{2}$. The invariance of Frobenius norm under unitary transformations reflects the profound connections between modulus and invariance in physics.
Remark 2.
In Theorem 2, $g_{\text{nonlin}}(\cdot)$ can handle equivariant features of any degree and the returned feature $\mathbf{v}$ has the same degree $l$ as the input feature $\mathbf{f}$. $\mathbf{v}$ is endowed with non-linear expressiveness with the effect of $s_{nonlin}(\cdot)$, which can be any differentiable non-linear network module, including non-linear activation function as a part. The gradient operation $\frac{\partial z}{\partial\mathbf{f}}$ in $g_{\text{nonlin}}(\cdot)$ is used during the forward inference process to promote feature encoding, expanding the usage and significance of gradients within the deep learning paradigm.
Taking $z$ in Eq. (5) as the bridge connecting Theorem 1 and Theorem 2, we present a general method for learning non-linear representations that are equivariant to SO(3), as shown in Fig. 2. The core of our method can be abstractly referred to as TraceGrad. At the label level, it incorporates the SO(3)-invariant trace quantity $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ introduced in Theorem 1 as a supervisory signal for learning SO(3)-invariant non-linear feature $z$. Given the attribute that invariance is preserved under non-linear operations, $z$ is encoded by $s_{nonlin}(\cdot)$ to obtain non-linear expressive capabilities. At the representation level, by taking the gradient of $z$ with respect to $\mathbf{f}$, the non-linear expressiveness of $z$ is transferred to the equivariant feature $\mathbf{v}$, while maintaining strict SO(3)-equivariance as we prove. Subsequently, merging $\mathbf{v}$ and $\mathbf{f}$, and applying $g_{\text{nonlin}}(\cdot)$ in a stacked manner can yield rich SO(3)-equivariant non-linear representations to predict the SO(3)-equivariant regression target. For details of the method, please refer to Section 6.
Figure 2: Methodological framework for learning SO(3)-equivariant representations with non-linear expressiveness to regress complex SO(3)-equivariant targets.
4 Results and Analyses
We apply our theory and method to the electronic-structure Hamiltonian prediction task, and collect results on eight benchmark databases, i.e., Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), Bilayer Bi2Se3 ($BS$), QH9-Stable (QS), and QH9-Dynamic (QD). The first six databases, consisting of crystal systems with elements like C, Mo, S, Bi, Te and Se, are from the DeepH benchmark series [6, 8]; while the last two databases, composed of molecular systems of C, H, O, N, F, are from the QH9 benchmark series [19] . These databases present diverse and complex challenges to the generality of a regression model. Regarding $MG$, $MM$, and QD, as their samples are prepared from a 300K environment, the thermal motions lead to complex non-rigid deformations, increasing the difficulty of Hamiltonian prediction. For $BG$, $BB$, $BT$, and $BS$, their twisted samples with an interplay of SO(3)-equivariant effects and van der Waals (vdW) force variations bring significant generalization challenges, which are further exacerbated by the absence of any twisted samples in the training sets. Besides, $BB$, $BT$, and $BS$ exhibit strong spin-orbit coupling (SOC) effects, which further increase the complexity of Hamiltonian modeling. For QS and QD, as we select the most difficult subset split strategies provided in their paper, namely ’ood’ and ’mol’, respectively, yielding differences in atom numbers and trajectories between training and testing samples, challenges in scale and trajectory extrapolation become prominent. Detailed information of these databases, including their statistics and visualizations, can be found in the Supplementary Information C.
We use a comprehensive set of metrics to deeply evaluate the performance of electronic-structure Hamiltonian prediction models. On the databases from the DeepH benchmark series [6, 8], we adpot the metrics of Mean Absolute Error (MAE) between predicted and ground truth Hamiltonians introduced by Yin et al. [17], including $MAE^{H}_{all}$ for measuring average MAE of all samples and matrix elements, $MAE^{H}_{cha\_s}$ for measuring the MAE of challenging samples where the baseline model performs the worst, $MAE^{H}_{block}$ for measuring the MAE of different basic blocks in the Hamiltonian matrix, and $MAE^{H}_{cha\_b}$ for measuring the MAE on the most challenging Hamiltonian block where the baseline model shows the poorest performance (with the largest $MAE^{H}_{block}$). These metrics comprehensively reflect the accuracy performance, covering not only the average accuracy but also the accuracy on difficult samples and challenging blocks of the Hamiltonian matrices. On the two databases from the QH9 benchmark series, we adopt the metrics introduced by their original paper [19], including MAE of Hamiltonian matrices, which are further subdivided into $MAE^{H}_{all}$ for measuring average MAE, $MAE^{H}_{diag}$ for measuring MAE of Hamiltonian matrix formed by an atom with itself, and $MAE^{H}_{non\_diag}$ for measuring MAE of Hamiltonian matrices formed by different atoms; as well as the MAE ($MAE^{\epsilon}$) of occupied orbital energies $\epsilon$ induced by the predicted Hamiltonians and compared to the ground truth ones, and the cosine similarity ($Sim(\psi)$) between the electronic wavefunctions $\psi$ induced by the predicted and ground truth Hamiltonians. $\epsilon$ and $\psi$ are crucial downstream physical quantities for determining multiple properties of the atomic systems as well as their dynamics, highly reflecting the application values of the Hamiltonian regression model.
We compare experimental results from two setups: the first one is the experimental results of the baseline SO(3)-equivariant regression model [8, 7] for Hamiltonian prediction, and the second one is the experimental results of extending the architecture and pipeline of the baseline model through the proposed TraceGrad method, which incorporates non-linear expressiveness into the SO(3)-equivariant features of the baseline model with the gradient operations of SO(3)-invariant non-linear features learned by the trace targets. We choose DeepH-E3 [8] as the baseline model for databases from the DeepH benchmark series [6, 8]; and we choose QHNet [7] as the baseline model for databases from the QH9 benchmark series [19]. The results of DeepH-E3 to be compared are taken from the reproduction of Yin et al. [17]. The reproduction, based on DeepH-E3’s newest open-source project [^1] https://github.com/Xiaoxun-Gong/DeepH-E3, showed slightly better results than that in the original publication, and was evaluated more comprehensively across multiple metrics. The results of QHNet to be compared are taken from their original paper [19], and for the unification of MAE units, we convert the units of MAE from $10^{-6}$ Hartree ($E_{h}$) in the original paper to meV222 $27211.41E_{h}=$ 27211.4 meV. The results of our methods are the average from 10 independent repeated experiments. To ensure reproducibility, we used a fixed random seed for all random processes. As a result, the standard deviation of the Hamiltonian prediction MAE across repeated experiments does not exceed $0.007$ meV, which is negligible.
We list the results of DeepH-E3 and DeepH-E3+TraceGrad in Tables 1 and 2 for monolayer structure databases ($MG$, $MM$) and bilayer structure databases ($BG$, $BB$, $BT$, $BS$) from the DeepH benchmark series, reporting the values of $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$, and $MAE^{H}_{cha\_b}$. Regarding the metric of $MAE^{H}_{block}$ for every Hamiltonian block, due to its large data volume, we just present its values for representative datasets here in Fig. 3, and attach its values for all databases from the DeepH series in Supplementary Information E. From these results, we could find that the proposed TraceGrad method dramatically enhances the accuracy performance of the baseline method DeepH-E3, both on average and for challenging samples or blocks, both on the non-twisted samples and the twisted samples. Specifically, on the corresponding datasets, TraceGrad lowers down the $MAE^{H}_{all}$ and $MAE^{H}_{cha}$ of DeepH-E3 with relative ratios of up to $34\%$ and $35\%$, respectively, and achieving better results in the vast majority of basic blocks. Particularly, for the blocks where DeepH-E3 perform the worst, TraceGrad reduces the $MAE$ ($MAE^{H}_{cha\_b}$) by a maximum of $48\%$. The leading performance on the $MG$ and $MM$ databases prepared at 300K demonstrates the robustness of our method against thermal motion. The high accuracy on the $BB$, $BT$, and $BS$ databases, which have strong SOC effects, indicates our method’s strong capability to model such effects. The excellent performance on the $BG^{t}$, $BB^{t}$, $BT^{t}$, and $BS^{t}$ subsets showcases the method’s superior generalization to twisted structures, which are not present in the training data. The outstanding performance on such samples highlights the good potential for studying twist-related phenomena, given interlayer twists are hot research topics because they may bring new electrical and transport properties [20, 21, 22]. Additionally, the $BG^{t}$ and $BB^{t}$ testing subsets contain significantly larger atomic systems compared to the training set (see Supplementary Information C for statistics of their sizes), yet our method still excels on these subsets for all metrics, demonstrating its good scalability on the sizes of atomic systems it handles.
We list the results of QHNet and QHNet+TraceGrad in Tables 3 and 4, for the $QS$ and $QD$ databases, respectively, reporting the values of $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$. From these results, we can observe that the proposed TraceGrad method significantly enhances the accuracy performance of the baseline method QHNet across all of these metrics on both of the two databases. Specifically, TraceGrad lowers down the $MAE^{H}_{all}$, $MAE^{H}_{non\_diag}$, $MAE^{H}_{diag}$, $MAE^{\epsilon}$ and $Sim(\psi)$ of QHNet with relative ratios of up to 40%, 39%, 40%, 76%, and 17%, on the corresponding databases. The significant improvements on the $QS$ database, partitioned using the ’ood’ split strategy [19] without scales overlapping among the training, validation, and testing sets, once again demonstrate the method’s strong generalization capabilities across different scales. Meanwhile, the improvements on the $QD$ database under the ’mol’ strategy [19] partitioning the training, validation, and testing sets with samples from totally different thermal motion trajectories, highlight the method’s robustness in generalizing to new thermal motion sequences. Furthermore, the significant promotion on the prediction accuracy of $\epsilon$, i.e., occupied orbital energies which are crucial for determining electronic properties such as optical characteristics and conductivity in atomic systems, and $\psi$, i.e., the electronic wavefunctions essential for understanding electron distribution and interactions, underscore good application potential of our method for tasks like material design and quantum computing.
The experimental results above indicate that the proposed representation learning framework, while maintaining strict SO(3)-equivariance, significantly enhances the expressiveness of equivariant representations, thus excellently capturing the complexity of physical systems and substantially breakthroughs the SOTA accuracy on the eight benchmark databases. The consistent improvements brought by our framework to different baseline models (DeepH-E3 or QHNet) and in various challenging scenarios, such as thermal motions, bilayer twists, scale variations, and new trajectories, fully demonstrate the universality and significance of our proposed theory and method.
Table 1: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$, and $MAE^{H}_{cha\_b}$ metrics (meV) on the Monolayer Graphene ($MG$) and Monolayer MoS2 ($MM$) databases. $\downarrow$ means lower values of the metrics correspond to better accuracy.
Methods $MG$ $MM$
$MAE$ ()($\downarrow$)( ↓ )
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.251 0.357 0.362 0.406 0.574 1.103
DeepH-E3+TraceGrad 0.175 0.257 0.228 0.285 0.412 0.808
Table 2: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$ and $MAE^{H}_{cha\_b}$ metrics (meV) on the Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$) and Bilayer Bi2Se3 ($BS$) databases, where the superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets. $\downarrow$ means lower values of the metrics correspond to better accuracy.
Methods $BG^{nt}$ $BG^{t}$
$MAE$ ()($\downarrow$)( ↓ )
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.389 0.453 0.644 0.264 0.429 0.609
DeepH-E3+TraceGrad 0.291 0.323 0.430 0.198 0.372 0.406
Methods $BB^{nt}$ $BB^{t}$
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.274 0.304 1.042 0.468 0.602 2.399
DeepH-E3+TraceGrad 0.226 0.256 0.740 0.384 0.503 1.284
Methods $BT^{nt}$ $BT^{t}$
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.447 0.480 1.387 0.831 0.850 4.572
DeepH-E3+TraceGrad 0.295 0.312 0.718 0.735 0.755 4.418
Methods $BS^{nt}$ $BS^{t}$
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.397 0.424 0.867 0.370 0.390 0.875
DeepH-E3+TraceGrad 0.300 0.332 0.644 0.291 0.302 0.674
Figure 3: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for representative datasets, where the superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets. Lower value of $MAE^{H}_{block}$ corresponds to better accuracy. The $MAE^{H}_{block}$ for other datasets from DeepH series can be found in Supplementary Information E.
Table 3: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ metrics on the QH9-Stable (QS) database using the ’ood’ split strategy [19]. $\downarrow$ means lower values of the metrics correspond to better accuracy, while $\uparrow$ means higher value of the metric corresponds to better performance. The units of MAE metrics are meV, while $Sim(\psi)$ is the cosine similarity which is dimensionless.
Methods MAE ()($\downarrow$)( ↓ ) $Sim(\psi)$ ()($\uparrow$)( ↑ )
$MAE^{H}_{all}$ $MAE^{H}_{diag}$ $MAE^{H}_{non\_diag}$ $MAE^{\epsilon}$
QHNet 1.962 3.040 1.902 17.528 0.937
QHNet+TraceGrad 1.191 2.125 1.139 8.579 0.948
Table 4: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ metrics on the QH9-Dynamic (QD) database using the ’mol’ split strategy [19]. $\downarrow$ means lower values of the metrics correspond to better accuracy, while $\uparrow$ means higher value of the metric corresponds to better performance. The units of MAE metrics are meV, while $Sim(\psi)$ is the cosine similarity which is dimensionless.
Methods MAE ()($\downarrow$)( ↓ ) $Sim(\psi)$ ()($\uparrow$)( ↑ )
$MAE^{H}_{all}$ $MAE^{H}_{diag}$ $MAE^{H}_{non\_diag}$ $MAE^{\epsilon}$
QHNet 4.733 11.347 4.182 264.483 0.792
QHNet+TraceGrad 2.819 6.844 2.497 63.375 0.927
We present experimental results of ablation study in Supplementary Information F. These fine-grained experiments assess the individual and combined contributions of the two core mechanisms of the proposed TraceGrad method on the six databases from the DeepH series. These mechanisms are the SO(3)-invariant trace supervision mechanism (Trace) at the label level and the gradient-based induction mechanism (Grad) at the representation layer. Results indicate that each mechanism can contribute individually to the performance. Moreover, their combination provides even better performance. This is because, on one hand, with the gradient-based induction mechanism as a bridge, the SO(3)-invariant features learned from the trace label can be transformed into the SO(3)-equivariant representations during inference; on the other hand, with trace label, the SO(3)-invariant network branch has a strong supervisory signal, enabling it to learn the intrinsic symmetry and complexity of the regression targets, enhancing the quality of SO(3)-invariant features and ultimately benefits the encoding of SO(3)-equivariant features. The value of such complementarity has been fully demonstrated in the experimental results.
5 Future Work
In future research, various extensions are conceivable across theoretical, methodological, and application fields:
First, from a theoretical perspective, we can generalize SO(3) group symmetry to a broader range and more complex groups. This will enable deep learning frameworks to accommodate richer mathematical structures.
Second, from a methodological perspective, we plan to combine our framework with advanced Transformer [23] and KAN [24] architectures to continuously enhance the expressive capabilities of equivariant features.
Third, from the application perspective, the current methodology extends beyond predicting electronic-structure Hamiltonians and is capable of generally predicting multiple physical quantities and properties that require equivariance. Furthermore, the methodology is not confined to physical research alone and may find applications in areas such as robotics, autonomous vehicles, and motion tracking systems. For instance, in the vision tasks of autonomous vehicles [25], where there is a high demand for robustness against coordinate transformations of cameras, the mainstream approach has been to achieve approximate SO(3)-equivariance through data augmentation. However, this does not guarantee absolute reliability and safety. Our work holds great potential to construct deep models with high generalization performance under the premise of strict SO(3)-equivariance, potentially revolutionizing techniques in such tasks. Looking forward to future endeavors, the foundational and pivotal role of this work is clear.
6 Method
6.1 Encoding and Decoding Framework
As shown in Fig. 2, our encoding framework corresponds to a total of $K$ encoding modules, which are sequentially connected to form a cascading stack. For the $k$-th module ($1\leq k\leq K$), it first introduces an SO(3)-equivariant base encoder, e.g., the encoder from DeepH-E3 [8] or QHNet [7] which is composed of recently developed equivariant operators [26, 27] like linear scaling, direct products, direct sums, Clebsch-Gordan decomposition, gated activation, equivariant normalization, and etc., to encode the physical system’s initial representations, such as spherical harmonics [28], or representations passed from the previous neural layers, as equivariant feature $\textbf{f}^{(k)}$ in the current hidden layer. Next, we construct the feature $\textbf{v}^{(k)}=g_{\text{nonlin}}(\textbf{f}^{(k)})$, to achieve sufficient non-linear expressiveness while maintaining SO(3)-equivariance, where $g_{\text{nonlin}}(\cdot)$ is the function defined in Eq. (5). $s_{nonlin}(\cdot)$, i.e., $g_{\text{nonlin}}(\cdot)$’s built-in non-linear mechanism, which can be any differentiable non-linear neural module, is implemented here as a feed-forward module with a hidden layer size of $1024$, using Layer Normalization for stabilization and regularization, and employing SiLU as the non-linear mapping. To facilitate learning and optimization, we connect the input and output of $g_{\text{nonlin}}(\cdot)$ in a residual manner [29], serving as the output of the current encoding module, i.e., $\textbf{o}^{(k)}=\textbf{f}^{(k)}+\textbf{v}^{(k)}$. $\textbf{o}^{(k)}$ is then fed into the next encoder module to yield $\textbf{f}^{(k+1)}$, $\textbf{o}^{(k+1)}$, …., until the last encoder module, which is followed by the decoder modules.
For an expressive feature encoder handling a complex physical system, $\textbf{f}^{(k)}$ is usually not a feature of single degree but a direct-sum concatenation of series of components $\textbf{f}^{(k)l_{1}}$, $\textbf{f}^{(k)l_{2}}$, …, $\textbf{f}^{(k)l_{k}}$ at multiple degrees, i.e., $l_{1}$, $l_{2}$, …, $l_{k}$, where some components share the same degree while others differ. In this case, it becomes necessary to extend the decomposition operator, i.e., $\text{CGDecomp}(\cdot)$ in Eq. (5), to accommodate various components of degrees. Moreover, in the context of neural networks, introducing learnable parameters $W$ may improve the model capacity. Based on these considerations, when constructing the encoding module, the $\text{CGDecomp}(\cdot)$ operation in Eq. (5) can be expanded as $\text{CGDecomp}_{ext}(\cdot)$:
$u^{(k)}=\text{CGDecomp}_{\text{ext}}(\textbf{f}^{(k)}\otimes\textbf{f}^{(k)},0,W)=\sum_{1\leq i,j\leq k,l_{i}=l_{j}}W_{ij}\cdot\text{CGDecomp}(\textbf{f}^{(k)l_{i}}\otimes\textbf{f}^{(k)l_{j}},0)$ (6)
where $W$ represents learnable parameters. This expansion does not change the equivariance or invariance of the features. It is obvious that, the output ($u^{(k)}$) of $\text{CGDecomp}_{\text{ext}}(\cdot)$ is also SO(3)-invariant, just like the original $\text{CGDecomp}(\cdot)$. Thus, even with such an expansion, the feature $z^{(k)}=s_{\text{nonlin}}(u^{(k)})$ remains SO(3)-invariant and $\textbf{v}^{(k)}=\frac{\partial z^{(k)}}{\partial\mathbf{f}^{(k)}}$ remain SO(3)-equivariant.
[…]
Since complex non-linear transformations have already been incorporated in the encoding framework, the decoders can be simple: the SO(3)-equivariant decoder is implemented using an SO(3)-equivariant linear mapping [27], and the SO(3)-invariant decoder is implemented as a shallow feed-forward module.
6.2 Training and Inference
The training loss function is shown as the following:
$\begin{split}\min_{\theta}loss&=loss_{Q}+\mu(loss_{Q},loss_{T})\cdot loss_{T},\\ loss_{Q}&=MSE(\widehat{\mathbf{Q}},\mathbf{Q}^{*}),\\ loss_{T}&=MSE(\widehat{\mathbf{T}},\mathbf{T}^{*})\end{split}$ (7)
where $\theta$ denotes all of the learnable parameters of our framework, $\widehat{\mathbf{Q}}$, $\widehat{\mathbf{T}}$ and $\mathbf{Q}^{*}$, $\mathbf{T}^{*}$ respectively denote the predictions and labels of $\mathbf{Q}$ and $\mathbf{T}$. In order that the both of the SO(3)-equivariant and SO(3)-invariant branches can be stably converged, we introduce $\mu(loss_{Q},loss_{T})$, i.e., a function returning a coefficient to regularize the relative scale of the two loss terms:
$\mu(loss_{Q},loss_{T})=\lambda\cdot No\_Grad(\frac{loss_{Q}}{loss_{T}})$ (8)
where $No\_Grad(\cdot)$ denotes gradient discarding when calculating such coefficient, as this coefficient is only used for adjusting the relative scale of the two loss terms and should not itself be a source of training gradients, otherwise it would counteract $loss_{T}$ in Eq. (7). The introduction of $\mu(loss_{Q},loss_{T})$ effectively prevents the numerical disparity between the two loss terms to stabilize the training for both of them. All of the encoding and decoding modules are trained jointly by Eq. (7), and the training configurations are included in Supplementary Information D.
In our framework, the SO(3)-invariant non-linear features $z^{(k)}(1\leq k\leq K)$ learned with the help of $\mathbf{T}$ contribute to the learning of SO(3)-equivariant representations $\textbf{v}^{(k)}(1\leq k\leq K)$ and the prediction of $\mathbf{Q}$. To achieve this goal, in the training phase, both the SO(3)-invariant feature encoder and the SO(3)-invariant decoder, used for predicting $\mathbf{T}$ from $z^{(k)}$, need to be trained. While in the inference phase, since $\mathbf{T}$ is not the final regression target, the SO(3)-invariant decoder can be removed, yet the SO(3)-invariant encoder still should be retained, used to induce SO(3)-equivariant features.
6.3 Time Complexity
Typically, one of the key advantage of deep learning algorithms for electronic-structure calculation lies in their low computational complexity which can asymptotically scale as $O(N)$, where $N$ is the number of atoms of a system. This is a significant efficiency advantage compared to the $O(N^{3})$ computational complexity of traditional DFT method, enabling efficient and low-cost prediction of electronic structures for large-scale atomic systems. Our method improves the accuracy of Hamiltonian predictions within the deep learning paradigm while maintaining this advantage of asymptotic complexity. The baseline models we select, whether DeepH-E3 or QHNet, are efficient models with an asymptotic computational complexity of $O(N)$ as $N$ increases. The neural module $g_{nonlin}(\cdot)$ we introduce efficiently operates on each individual feature f from the baseline models to obtain the corresponding non-linear SO(3)-equivariant feature v, rather than relying on complex interactions between multiple features to derive new ones. This ensures that the computational complexity remains asymptotically $O(N)$, with good potential to efficiently handle large atomic systems.
7 Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant Number 12134012) and the Strategic Priority Research Program of Chinese Academy of Sciences (Grant Number XDB0500201).
Appendix A The Equivariance-Expressiveness Dilemma in Existing Deep Learning Methods
In this section, we introduce existing efforts to harmonize the deep learning paradigm with SO(3)-equivariance, analyzing the dilemma between strict SO(3)-equivariance and non-linear expressive capabilities. This dilemma leads to an unsatisfactory trade-off between strict equivariance and non-linear expressiveness, ultimately limiting the achievable performance and scopes of application for these regression models.
A.1 Approaches with Strict SO(3)-equivariance but Weak Expressiveness
Most existing methods have focused on achieving SO(3)-equivariance at the neural mechanism level by applying group theory-guaranteed SO(3)-equivariant operators [26, 27]. These operators, including linear scaling, element-wise sum, direct products, direct sums, and Clebsch-Gordan decomposition, have been used to build graph neural network architectures for tasks in 3D point cloud analysis [30], molecular dynamic simulation and property prediction [13, 12], as well as Hamiltonian prediction [31, 4, 8, 14]. However, these operators could not provide sufficient expressiveness to fit complex non-linear mappings. The majority of these operators are linear operations, and the only one (direct product) that introduces some non-linearity can only approximate a simple squaring function, and cannot be extended to more complex non-linear forms due to the huge computational burden and difficulty in convergence. Even worse, directly applying highly-expressive non-linear activation functions, such as $Sigmoid$, $Softmax$, and $SiLU$, to SO(3)-equivariant features may result in the loss of equivariance. To alleviate this issue, these methods introduced a gated activation mechanism that processes SO(3)-invariant features ($l=0$) using non-linear activations and employs them as linear gating coefficients for SO(3)-equivariant features ($l\geq 1$), aiming to enhance their expressive power while maintaining strict equivariance. However, from the perspective of SO(3)-equivariant features, this mechanism is a combination of linear operations and cannot fundamentally increase their non-linear expressive capabilities.
A.2 Approaches with Remarkable Expressiveness but Approximate SO(3)-equivariance
Li et al. [6] proposed a local coordinate strategy to achieve SO(3)-equivariance. This strategy relied on the transformation relationship between local and global coordinates to achieve equivariance, without imposing restrictions on the neural network, allowing the use of network modules that possess non-linear expressiveness. However, this strategy, solely depending on coordinate transformations, only works under holistic rigid rotations of the atomic system. Without exploration of a neural mechanism for equivariance, when the local coordinates experience non-rigid perturbations such as thermal movements or bilayer twists, this strategy offers no guarantees on maintaining symmetry properties; Zitnick et al. [11] decomposed features into weighted sums of spherical harmonic basis functions, each corresponding to a discrete point on a unit sphere, through inner-product operations. Such discrete decomposition enables non-linear operations on the coefficients to enhance expressiveness. In subsequent developments [15, 16, 32], this approach has demonstrated a remarkable capacity to fit complex functions. However, the inner-product operations with discrete basis make such approach degenerate from continuous to discrete rotational equivariance, no longer maintaining strict equivariance to continuous rotational operations, as pointed out by Zhang et al. [1]. Additionally, due to the high complexity of its discrete sampling strategies, tensor orders ($m$) corresponding to large degrees ($l$) are often partially pruned in architectures such as Equiformerv2 [16], further sacrificing the equivariance of features. In summary, although these methods with approximate SO(3)-equivariance achieve high numerical accuracy due to their extensive use of non-linear mechanisms, they are best suited for the applications that primarily require numerical approximations but not for tasks that demand theoretical depth in physics. In some applications, the demands for symmetry are extremely high. Even minor deviations from perfect symmetry can result in incorrect physical results.
Combining these analyses, existing deep learning methods have compromised either strict SO(3)-equivariance or non-linear expressiveness when processing physical systems. In this paper, we develop a novel framework that theoretically combines strict SO(3)-equivariance with the non-linear expressiveness of neural networks. This solution resolves the equivariance-expressiveness dilemma by deeply exploring the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities and their representations.
Appendix B Proof of Theory
B.1 Proof of Theorem 1
Proof.
Under an SO(3) rotation represented by the rotational matrix $\mathbf{R}$, $\mathbf{Q}=\mathbf{Q}^{l_{p}\otimes l_{q}}$ is transformed as $\mathbf{Q}(\mathbf{R})$:
$\mathbf{Q}(\mathbf{R})=\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger},$
where $\mathbf{D}^{l_{p}}(\mathbf{R})$ and $\mathbf{D}^{l_{q}}(\mathbf{R})$ are the Wigner-D matrices for the degrees of $l_{p}$ and $l_{q}$, respectively, corresponding to the rotation $\mathbf{R}$.
The conjugate transpose of the transformed quantity is:
$\mathbf{Q}(\mathbf{R})^{\dagger}=\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}.$
Thus, the product $\mathbf{Q}(\mathbf{R})\cdot\mathbf{Q}(\mathbf{R})^{\dagger}$ is:
$\mathbf{Q}(\mathbf{R})\cdot\mathbf{Q}(\mathbf{R})^{\dagger}=(\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger})\cdot(\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}).$
Using the cyclic property of the trace, which states that the trace of a product of matrices remains unchanged under cyclic permutations (i.e., $tr(ABC)=tr(BCA)=tr(CAB)$), and combining the properties that $\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}=\mathbf{I}$ and $\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger}=\mathbf{I}$, we can rearrange the terms inside the trace as follows:
$\mathbf{T}(\mathbf{R})=tr(\mathbf{Q}(\mathbf{R})\cdot\mathbf{Q}(\mathbf{R})^{\dagger})=tr((\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger})\cdot(\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}))$
$=tr(\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger})=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R}))=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\mathbf{T}.$
Therefore, $\mathbf{T}=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is invariant under SO(3) transformations, its SO(3)-invariance is proved. ∎
B.2 Proof of Theorem 2
Proof.
The input feature $\mathbf{f}$ is SO(3)-equivariant, meaning that under an SO(3) rotation represented by $\mathbf{R}$, it transforms as follows:
$\mathbf{f}(\mathbf{R})=\mathbf{D}^{l}(\mathbf{R})\cdot\mathbf{f}$
where $\mathbf{D}^{l}(\mathbf{R})$ is the Wigner-D matrix corresponding to degree $l$.
First, according to group theory, $u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ is an SO(3)-invariant scalar because the degree-zero component from the Clebsch-Gordan decomposition is invariant under rotations. Since applying a non-linear operation to an SO(3)-invariant quantity does not change its invariance, $z=s_{\text{nonlin}}(u)$ is also SO(3)-invariant, independent to the specific form of $s_{\text{nonlin}}(\cdot)$. It formally holds that:
$z=z(\mathbf{R})$ (9)
Next, we apply the chain rule in Jacobian form. Considering $\mathbf{f}(\mathbf{R})$ is a column vector, to facilitate the application of the chain rule in vector form, we first transpose it into a row vector $\mathbf{f}(\mathbf{R})^{T}$, then differentiate:
$\frac{\partial z(\mathbf{R})}{\partial\mathbf{f}^{T}(\mathbf{R})}=\frac{\partial z}{\partial\mathbf{f}^{T}(\mathbf{R})}=\frac{\partial z}{\partial\mathbf{f}^{T}}\frac{\partial\mathbf{f}^{T}}{\partial\mathbf{f}(\mathbf{R})^{T}}=\frac{\partial z}{\partial\mathbf{f}^{T}}\cdot\mathbf{D}^{l}(\mathbf{R})^{-1}=\frac{\partial z}{\partial\mathbf{f}^{T}}\cdot\mathbf{D}^{l}(\mathbf{R})^{T}$ (10)
Here we utilize the property that $\mathbf{D}^{l}(\mathbf{R})^{-1}=\mathbf{D}^{l}(\mathbf{R})^{T}$. Since the representations of neural networks are generally real numbers, the Wigner-D matrix is also real unitary  [^3] In Theorem 1 and Theorem 2, the Wigner-D matrices are in the complex and real fields, respectively, because the output space may be complex while the internal representations of neural networks are typically real. Although the internal representations of neural network modules including $g_{\text{nonlin}}(\cdot)$ are in the real field, they are still able to predict targets in complex field with equivariance. Literature [8] has provided mechanisms for converting the network outputs in the real field into regression targets with real and imaginary parts..
Finally, we transpose the result back to a column vector:
$\mathbf{v}(\mathbf{R})=g_{nonlin}(\mathbf{f}(\mathbf{R}))=(\frac{\partial z(\mathbf{R})}{\partial\mathbf{f}^{T}(\mathbf{R})})^{T}=(\frac{\partial z}{\partial\mathbf{f}^{T}}\cdot\mathbf{D}^{l}(\mathbf{R})^{T})^{T}=\mathbf{D}^{l}(\mathbf{R})\cdot\frac{\partial z}{\partial\mathbf{f}}=\mathbf{D}^{l}(\mathbf{R})\cdot\mathbf{v}$ (11)
This proves that $g_{\text{nonlin}}(\cdot)$ is an SO(3)-equivariant non-linear operator: when applying its non-linearity to a SO(3)-equivariant feature $\mathbf{f}$, the output feature $\mathbf{v}$ remains SO(3)-equivariant. ∎
Appendix C Information of Experimental Databases
In this part, we provide detailed information about the experimental databases, including the statistical information of the six databases from the DeepH series [6, 8], and the two databases from the QH9 series [19], listed in Table 5 and Table 6, respectively. Additionally, we visualize two types of challenging testing samples: samples with non-rigid deformation from thermal motions, as well as the bilayer samples with interlayer twists, which are shown in Fig. 4 and Fig. 5, respectively.
Table 5: Statistic information of the six benchmark databases, i.e., Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), Bilayer Bi2Se3 ($BS$), from the DeepH benchmark series [6, 8]. SOC: effects of Spin-Orbit Coupling. $m$: number of samples in the current dataset; $a_{max}$: maximum number of atoms for a sample in the current dataset. $a_{min}$: minimum number of atoms for a sample in the current dataset. $nt$: non-twisted samples. $t$: twisted samples.
Statistic Types MG MM BG BB BT BS
Elements C Mo, S C Bi Bi, Te Bi, Se
SOC weak weak weak strong strong strong
Training ($nt$) $m$ 270 300 180 231 204 231
$a_{max}$ 72 75 64 36 90 90
$a_{min}$ 72 75 64 36 90 90
Validation ($nt$) $m$ 90 100 60 113 38 113
$a_{max}$ 72 75 64 36 90 90
$a_{min}$ 72 75 64 36 90 90
Testing ($nt$) $m$ 90 100 60 113 12 113
$a_{max}$ 72 75 64 36 90 90
$a_{min}$ 72 75 64 36 90 90
Testing ($t$) $m$ 9 4 2 2
$a_{max}$ 1084 244 130 190
$a_{min}$ 28 28 70 70
Table 6: Basic statistics of the two benchmark databases, QH9-Stable ($QS$) and QH9-Dynamic ($QD$), from the QH9 benchmark series [19]. For the $QS$ database, the ’ood’ strategy from the official settings is used to split the training, validation, and test sets, ensuring that the atom number of samples do not overlap across the three subsets. For the $QD$ database, the ’mol’ strategy from the official settings is employed to split the training, validation, and test sets, ensuring that there are no thermal motion samples from the same temporal trajectory across the three subsets. These split strategies aim to assess the regression model’s extrapolation capability with respect to the number of atoms as well as the temporal trajectories. $m$: number of samples in the current dataset. $a_{max}$: maximum number of atoms for a sample in the current dataset. $a_{min}$: minimum number of atoms for a sample in the current dataset.
Statistic Types QS QD
Elements C, H, O, N, F C, H, O, N, F
Training $m$ 104,001 79,900
$a_{max}$ 20 19
$a_{min}$ 3 10
Validation $m$ 17,495 9,900
$a_{max}$ 22 19
$a_{min}$ 21 10
Testing $m$ 9,335 10,100
$a_{max}$ 29 19
$a_{min}$ 23 10
Figure 4: Visualization of testing samples exhibiting non-rigid deformations due to thermal motions
Figure 5: Visualization of testing samples with interlayer twists.
Appendix D Training Configurations
$\lambda$ in the training loss function is set as $0.3$ according to parameter selection on the validation sets. The initial learning rates for experiments on the $MG$, $MM$, $BG$, $BB$, $BT$ and $BS$ databases are set as $0.003$, $0.005$, $0.003$, $0.005$, $0.004$, and $0.005$, respectively, the training batch size for them is set as $1$ and the optimizer is selected as Adam; the initial learning rate for experiments on the $QS$ and $QD$ databases is set as $0.0005$, the training batch size is set as $16$, and the optimizer is selected as AdamW.
Appendix E Visualization of Block-level MAE Statistics
The MAE metrics (denoted as $MAE^{H}_{block}$) for the two setups, i.e., DeepH-E3 and DeepH-E3+TraceGrad, across different blocks of the Hamiltonian matrix for six databases from the DeepH benchmark series are illustrated in Fig. 6 and 7. Fig. 6 presents the results for monolayer structures, while Fig. 7 focuses on bilayer structures. From these figures, it can be observed that our method, TraceGrad, brings significant accuracy improvements over the baseline method, i.e., DeepH-E3, on most of the blocks of the Hamiltonian matrix, especially on those blocks where DeepH-E3 performs poorly.
Figure 6: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the $MG$ and $MM$ databases.
Figure 7: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the non-twisted (marked with superscripts $nt$) and twisted (marked with superscripts $t$) structures of $BG$, $BB$, $BT$ and $BS$.
Appendix F Ablation Study
We conduct fine-grained ablation study on the six databases from DeepH benchmark series, comparing results from the four setups:
• DeepH-E3 [8]: the baseline model.
• DeepH-E3+Trace: this experimental setup, an ablation term, only implements half part of our method. Specifically, it extends the architecture of DeepH-E3 by adding our SO(3)-invariant encoding and decoding branches and using the trace quantity $\mathbf{T}=\text{tr}(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\text{tr}(\mathbf{H}^{l_{p}\otimes l_{q}}\cdot(\mathbf{H}^{l_{p}\otimes l_{q}})^{\dagger})$ to train them. As for ablation study, this setup does not include the gradient-based mechanism for encoding SO(3)-equivariant non-linear features; instead, it directly uses the SO(3)-equivariant features outputted by DeepH-E3 for Hamiltonian regression.
• DeepH-E3+Grad: this setup is also an ablation term and implements the other half part of our method in contrast to the previous ablation term. Specifically, it incorporates our SO(3)-invariant encoder branch as well as the gradient-induced operator to yield SO(3)-equivariant non-linear features based on the representations provided by the baseline encoders of DeepH-E3 for Hamiltonian prediction. As for ablation study, this setup continues to use the single-task training pipeline of DeepH-E3, predicting only the Hamiltonian without joint supervision training through the trace label.
• DeepH-E3+TraceGrad: this is a complete implementation of our framework extending beyond of the architecture and training pipeline of DeepH-E3, at the label level, we introduce the trace quantity to guide the learning of SO(3)-invariant features; Meanwhile, at the representation level, we leverage the gradient operator to yield SO(3)-equivariant non-linear features for Hamiltonian prediction.
Table 7: Ablation study MAE results (meV) on the $MG$ and $MM$ databases. $\downarrow$ means lower values of the metrics correspond to better accuracy.
Methods $MG$ $MM$
$MAE$ ()($\downarrow$)( ↓ )
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.251 0.357 0.362 0.406 0.574 1.103
DeepH-E3+Trace 0.230 0.344 0.348 0.378 0.537 1.091
DeepH-E3+Grad 0.185 0.269 0.258 0.308 0.453 0.924
DeepH-E3+TraceGrad 0.175 0.257 0.228 0.285 0.412 0.808
Table 8: Ablation study MAE results (meV) on $BG$, $BB$, $BT$and $BS$. The superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets. $\downarrow$ means lower values of the metrics correspond to better accuracy.
Methods $BG^{nt}$ $BG^{t}$
$MAE$ ()($\downarrow$)( ↓ )
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.389 0.453 0.644 0.264 0.429 0.609
DeepH-E3+Trace 0.362 0.417 0.593 0.251 0.401 0.480
DeepH-E3+Grad 0.320 0.356 0.511 0.222 0.389 0.446
DeepH-E3+TraceGrad 0.291 0.323 0.430 0.198 0.372 0.406
Methods $BB^{nt}$ $BB^{t}$
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.274 0.304 1.042 0.468 0.602 2.399
DeepH-E3+Trace 0.259 0.285 0.928 0.429 0.570 1.782
DeepH-E3+Grad 0.243 0.272 0.824 0.406 0.542 1.431
DeepH-E3+TraceGrad 0.226 0.256 0.740 0.384 0.503 1.284
Methods $BT^{nt}$ $BT^{t}$
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.447 0.480 1.387 0.831 0.850 4.572
DeepH-E3+Trace 0.406 0.462 1.239 0.784 0.812 4.520
DeepH-E3+Grad 0.342 0.365 0.750 0.742 0.786 4.463
DeepH-E3+TraceGrad 0.295 0.312 0.718 0.735 0.755 4.418
Methods $BS^{nt}$ $BS^{t}$
$MAE_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.397 0.424 0.867 0.370 0.390 0.875
DeepH-E3+Trace 0.382 0.397 0.843 0.351 0.367 0.838
DeepH-E3+Grad 0.343 0.365 0.696 0.324 0.339 0.746
DeepH-E3+TraceGrad 0.300 0.332 0.644 0.291 0.302 0.674
Experimental results of the four setups are listed in Table 7 and 8. From the results of ablation terms, we can obtain a more fine-grained experimental analysis. By comparing among the results of DeepH-E3, DeepH-E3+Trace, DeepH-E3+Grad, and DeepH-E3+TraceGrad, we can conclude that the two core mechanisms of our method, i.e., the SO(3)-invariant trace supervision mechanism (Trace) at the label level as well as the gradient-based induction mechanism (Grad) at the representation layer, can contribute to the performance individually. Moreover, their combination provides even better performance.