### v4

# A Framework of SO(3)-equivariant Non-linear Representation Learning and its Application to Electronic-Structure Hamiltonian Prediction

Abstract
We propose both a theoretical and a methodological framework to address a critical challenge in applying deep learning to physical systems: the reconciliation of non-linear expressiveness with SO(3)-equivariance in predictions of SO(3)-equivariant quantities. Inspired by covariant theory in physics, we present a solution by exploring the mathematical relationships between SO(3)-invariant and SO(3)-equivariant quantities and their representations. We first construct theoretical SO(3)-invariant quantities derived from the SO(3)-equivariant regression targets, and use these invariant quantities as supervisory labels to guide the learning of high-quality SO(3)-invariant features. Given that SO(3)-invariance is preserved under non-linear operations, the encoding process for invariant features can extensively utilize non-linear mappings, thereby fully capturing the non-linear patterns inherent in physical systems. Building on this, we propose a gradient-based mechanism to induce SO(3)-equivariant encodings of various degrees from the learned SO(3)-invariant features. This mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations, while theoretically preserving their equivariant properties as we prove, establishing a strong foundation for regressing complex SO(3)-equivariant targets. We apply our theory and method to the electronic-structure Hamiltonian prediction tasks, experimental results on eight benchmark databases covering multiple types of systems and challenging scenarios show substantial improvements on the state-of-the-art prediction accuracy of deep learning paradigm. Our method boosts Hamiltonian prediction accuracy by up to 40% and enhances downstream physical quantities, such as occupied orbital energy, by a maximum of 76%. Our method also significantly promotes the acceleration ratios for the convergence of traditional Density Functional Theory (DFT) methods.
1 Introduction
With the advantages in computational complexity and generalization capabilities, deep learning paradigm has vigorously driven advancements in physics research (Zhang et al., 2023). For example, in predicting electronic-structure Hamiltonians, traditional Density Functional Theory (DFT) methods (Hohenberg & Kohn, 1964; Kohn & Sham, 1965) suffer from computational complexities of $\mathcal{O}(N^{3})$, where $N$ is the number of atoms in a system. In contrast, deep learning approaches have significantly reduced this complexity to $\mathcal{O}(N)$ (Unke et al., 2021; Gu et al., 2022; Li et al., 2022; Yu et al., 2023b; Gong et al., 2023), opening up new possibilities for analyzing extremely large atomic systems, enabling efficient materials simulation and design, as well as groundbreaking molecular pharmaceutical research that was previously unimaginable.
However, deep learning methods still face substantial challenges when processing physical systems. To align with fundamental physical laws, these methods must strictly adhere to symmetry principles. For example, physical quantities such as force fields and electronic-structure Hamiltonians must be equivariant under 3D rotational transformations, i.e., elements from the SO(3) group. Besides, the calculation of physical quantities calls for high numerical accuracy, necessitating that the neural networks possess strong capabilities to express the complex non-linear mappings from atomic structures to the regression targets. However, it is challenging for deep learning methods to simultaneously ensure strict SO(3)-equivariance and high numerical accuracy on modeling physical systems. The root cause of this problem lies in the conflict between SO(3)-equivariance and non-linear expressiveness: specifically, directly applying non-linear activation functions on SO(3)-equivariant features (with degree $l\geq 1$) may lead to the loss of equivariance, while bypassing non-linear mappings severely restricts the network’s expressive capabilities and thereby lowering down the achievable accuracy. This issue is commonly found in physics-oriented machine learning tasks that demand both strict equivariance and fine-grained generalization performance, as analyzed by Zitnick et al. (2022). Several recent efforts have been made to alleviate this issue (Zitnick et al., 2022; Passaro & Zitnick, 2023; Wang et al., 2024c; Yin et al., 2024), however, they often compromised strict SO(3)-equivariance.
To address the equivariance-expressiveness dilemma, we make theoretical and methodological explorations on unifying strict SO(3)-equivariance with strong non-linear expressiveness within the realm of deep representation learning. We are inspired by the insight that invariant quantities in transformation often reflect the mathematical nature of physical laws and can induce other quantities with equivariant properties. For example, in special relativity (Resnick, 1991), the spacetime interval between events is invariant under Lorentz transformations, and is fundamental in formulating physical laws that involve equivariant quantities. For another example, the total electronic energy of an atomic system is SO(3)-invariant, while the derived force fields are SO(3)-equivariant. We aim to extend the relationship between invariance and equivariance from specific physical quantities to a more general representation learning framework. From the perspective of deep representation learning, the attribute that invariance is preserved under non-linear operations is a significant advantage, given its compatibility with non-linear expressive capabilities. Built upon these insights, we propose a solution to the equivariance-expressiveness dilemma by intensively exploring and making use of the intrinsic relationships between SO(3)-invariant and SO(3)-equivariant quantities and representations: we first dedicate efforts to learning high-quality SO(3)-invariant features with ample non-linear expressiveness, and subsequently, we derive SO(3)-equivariant non-linear representations and the target quantities from these SO(3)-invariant ones. Specifically:
First, we propose a theoretical construct of SO(3)-invariant quantities, namely $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$, where $tr(\cdot)$ signifies the trace operation, $\dagger$ denotes the conjugate transpose operation, and $\mathbf{Q}$ denotes the SO(3)-equivariant regression targets we aim to predict. A significant advantage of these SO(3)-invariant quantities lies in the fact that they are directly derived from the SO(3)-equivariant target labels and can serve as unique supervision labels for the effective learning of informative SO(3)-invariant features that capture the intrinsic symmetry properties of the mathematical structure of $\mathbf{Q}$ without requiring additional labeling resources.
Second, we propose a gradient-based mechanism to induce SO(3)-equivariant representations for predicting the regression target $\mathbf{Q}$ in the inference phase from high-quality non-linear SO(3)-invariant features learned under the supervision of $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ in the training phase. Taking SO(3)-invariant features as a bridge, this mechanism can incorporate non-linear expressive capabilities into SO(3)-equivariant representations while preserving their equivariant properties, as we prove, laying a solid foundation for accurately inferring complex SO(3)-equivariant targets.
We develop our theory into a unified SO(3)-equivariant non-linear representation learning method, and apply it to the computation of electronic-structure Hamiltonian, which occupies a central position in quantum mechanics to induce key physical quantities such as electronic wavefuction, orbital energy, and band structure, bringing deep understanding of the electrical, optical, magnetic, and transport properties of atomic systems, while posing significant difficulties for machine learning techniques due to its intrinsic symmetry with respect to the SO(3) group as well as the high-dimensional complexity, as pointed out by Yin et al. (2024). Our method significantly improves the state-of-the-art performance in Hamiltonian prediction on eight databases from the well-known DeepH and QH9 benchmark series (Li et al., 2022; Gong et al., 2023; Yu et al., 2023a). It demonstrates excellent generalization performance to both crystalline and molecular systems, covering challenging scenarios such as thermal motions, bilayer twists, scale variations, and new trajectories. Furthermore, as observed from the experiments on the QH9 benchmark series, our approach also substantially enhances the prediction accuracy of downstream physical quantities of Hamiltonian including occupied orbital energy and electronic wavefunction. Moreover, our method also demonstrates superior performance in accelerating the convergence of classical DFT by providing predicted Hamiltonians as initialization matrices. Our leading performance comprehensively demonstrates that our method, while satisfying SO(3)-equivariance, possesses excellent expressive power and generalization performance, providing an effective deep learning tool for efficient and accurate electronic-structure calculations of atomic systems.
2 Related Work
The SO(3)-equivariant representation learning paradigm (Thomas et al., 2018; Geiger & Smidt, 2022) typically developed group theory-based symmetry operators, such as linear scaling, element-wise sum, direct products, direct sums, Clebsch-Gordan decomposition, and equivariant normalization, to encode equivariant features. These operators have been used to construct graph neural network architectures for tasks in 3D point cloud analysis (Fuchs et al., 2020), molecular property prediction and dynamic simulation (Musaelian et al., 2023; Liao & Smidt, 2023), as well as Hamiltonian prediction (Schütt et al., 2019; Unke et al., 2021; Gong et al., 2023; Zhong et al., 2023). However, as non-linear activation functions may result in the loss of equivariance, they are restricted when applied to SO(3)-equivariant features with degree $l$ greater than one. This restriction severely limits the network’s capability to model complex non-linear mappings. To alleviate this issue, methods like DeepH-E3 (Gong et al., 2023), QHNet (Yu et al., 2023b) and Equiformer (Liao & Smidt, 2023) introduced a gated activation mechanism that feeds SO(3)-invariant features ($l=0$) into non-linear activation functions and uses these features as linear gating coefficients for SO(3)-equivariant features ($l\geq 1$), aiming to enhance their expressive power while maintaining strict equivariance. Nevertheless, multiplying SO(3)-equivariant features with linear coefficients may not fundamentally improve their non-linear expressiveness.
In order to improve non-linear expressiveness, Zitnick et al. (2022) decomposed SO(3)-equivariant features into SO(3)-invariant coefficients of spherical harmonic basis functions. These SO(3)-invariant coefficients were processed by non-linear neural networks to enhance expressiveness, with equivariance regained by recombining the updated coefficients with the basis functions. In subsequent developments (Passaro & Zitnick, 2023; Liao et al., 2024; Wang et al., 2024b; c), this approach has demonstrated a remarkable capacity to fit complex functions. However, as pointed out by existing literature (Zhang et al., 2023), this approach degenerates from continuous to discrete rotational equivariance, losing strict equivariance to continuous rotational transformations due to the decomposition based on inner-product operations with discrete basis functions; Li et al. (2022) proposed a local coordinate strategy, projecting rotating global coordinates onto SO(3)-invariant local ones for the non-linear neural network to encode. However, this strategy is effective only for global rigid rotations and fails to maintain symmetry under non-rigid perturbations like thermal fluctuations or bilayer twists, as it lacks a neural mechanism to enforce equivariance; Yin et al. (2024) proposed a hybrid framework consisting of both group theory-guaranteed SO(3)-equivariant mechanisms and non-linear mechanisms to regress Hamiltonians. In this framework, the non-linear mechanisms showed remarkable capability at learning SO(3)-equivariance from the data with the help of the theoretically SO(3)-equivariant mechanisms, and released powerful non-linear expressive capabilities to achieve more numerical accuracy. However, the equivariance achieved through data-driven methods does not have strict theoretical guarantee, even with rotational data augmentation. In some applications, the demands for symmetry are extremely high. Even minor deviations from perfect SO(3)-equivariance can result in incorrect physical results. In this paper, we propose a framework that theoretically combines strict SO(3)-equivariance with the non-linear expressiveness of neural networks to resolve the equivariance-expressiveness dilemma.
3 Problem Formalization
For a background introduction on groups, group representations, equivariance, and invariance, please refer to Appendix A. Here, we directly focus on the equivariance of physical quantities under 3D rotational operations that form the SO(3) group. Let ${\bf{Q}}^{l_{p}\otimes l_{q}}$ denote an SO(3)-equivariant quantity in direct-product state formed by $l_{p}\otimes l_{q}$, i.e., the direct product between degrees $l_{p}$ and $l_{q}$. It obeys the following SO(3)-equivariant law:
${\bf{Q}}({\bf{R}})^{l_{p}\otimes l_{q}}={\bf{D}}^{l_{p}}({\bf{R}})\cdot{\bf{Q}}^{l_{p}\otimes l_{q}}\cdot{({\bf{D}}^{l_{q}}({\bf{R}}))^{\dagger}}$ (1)
where $\dagger$ denotes the conjugate transpose operation. ${\bf{R}}\in\mathbb{R}^{3\times 3}$ is the rotational matrix, ${\bf{D}}^{l_{p}}({\bf{R}})\in\mathbb{R}^{(2l_{p}+1)\times(2l_{p}+1)}$ and ${\bf{D}}^{l_{q}}({\bf{R}})\in\mathbb{R}^{(2l_{q}+1)\times(2l_{q}+1)}$ are the Wigner-D matrices of degrees $l_{p}$ and $l_{q}$, respectively; ${\bf{Q}}({\bf{R}})^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ denotes the transformed results of ${\bf{Q}}^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ through the rotational operation by ${\bf{R}}$.
${\bf{Q}}^{l_{p}\otimes l_{q}}$ in the direct-product state can be further decomposed into a series of direct-sum state components, i.e., ${\bf{q}}^{l}(|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q})$, which follows SO(3)-equivariant law mathematically equivalent to Eq. 1 but with a simpler form:
${\bf{q}}({\bf{R}})^{l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{q}}^{l},|l_{p}-l_{q}|\leq l\leq l_{p}+l_{q}$ (2)
where ${\bf{q}}^{l}\in\mathbb{R}^{2l+1}$ and ${\bf{q}}({\bf{R}})^{l}\in\mathbb{R}^{2l+1}$ respectively denote the components with degree $l$ before and after the rotational operation by ${\bf{R}}$.
For ease of processing, the internal representations of SO(3)-equivariant neural networks (Gong et al., 2023; Liao & Smidt, 2023) are typically in the direct-sum form. To obey the equivariant law of Eq. 2 for the regression target, these hidden representations must also satisfy the same form of equivariance:
${\bf{f}}({\bf{R}})^{(k)l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{f}}^{(k)l}$ (3)
where ${\bf{f}}^{(k)l}\in\mathbb{R}^{2l+1}$ and ${\bf{f}}({\bf{R}})^{(k)l}\in\mathbb{R}^{2l+1}$ respectively denote one channel of hidden features with degree $l$ before and after the rotational operation by ${\bf{R}}$, at the $k$ th hidden layer.
Due to the intrinsic complexity and non-linearity of physic quantities, neural networks on regressing these quantities are supposed to equip with non-linear mappings to fully capture the intrinsic patterns of the physical quantities, which is crucial for precise and generalizable prediction performance. Meanwhile, the non-linear mappings, denoted as $g_{nonlin}(\cdot)$, must also preserve SO(3)-equivariance, which is expressed as:
${\bf{f}}({\bf{R}})^{(k+1)l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{f}}^{(k+1)l},\,\,\,\,\text{subject to}\,\,{\bf{f}}^{(k+1)l}=g_{nonlin}({\bf{f}}^{(k)l})$ (4)
However, directly implementing $g_{nonlin}(\cdot)$ as neural network module with non-linear activation functions, such as $Sigmoid$, $Softmax$ and $SiLU$, may result in the destruction of strict equivariance. How to make $g_{nonlin}(\cdot)$ both theoretically SO(3)-equivariant and capable of non-linear expressiveness, and effectively apply it to the prediction of SO(3)-equivariant complex physical quantities, i.e., electronic-structure Hamiltonian in this context, is the core problem this paper aims at solving.
4 Theory
Theorem 1.
The quantity $\mathbf{T}=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is SO(3)-invariant, where $\mathbf{Q}$ is the simplified representation (without superscripts) of $\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 3, and $\dagger$ denotes the conjugate transpose operation, $tr(\cdot)$ is the trace operation.
Theorem 2.
The non-linear neural mapping $g_{nonlin}(\cdot)$ defined as the following is SO(3)-equivariant:
$\mathbf{v}=g_{nonlin}(\mathbf{f})=\frac{\partial z}{\partial\mathbf{f}},\quad\text{subject to}\quad z=s_{nonlin}(u),u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ (5)
where $\mathbf{f}$ is an input SO(3)-equivariant feature with degree $l$ in the direct-sum state, $\otimes$ denotes the direct-product operation of tensors, $\text{CGDecomp}(\cdot,0)$ refers to performing a Clebsch-Gordan decomposition of the tensor and returning the scalar component of degree 0, $s_{nonlin}(\cdot)$ represents arbitrary differentiable non-linear neural modules, and $\mathbf{v}$ is the outputted feature encoded by $g_{nonlin}(\cdot)$.
The proofs of the two theorems above are presented in Appendix B.
Figure 1: Methodological framework for learning SO(3)-equivariant representations with non-linear expressiveness to regress complex SO(3)-equivariant targets.
5 Method
As shown in Fig. 1, taking $z$ in Eq. 5 serving as the bridge between Theorem 1 and Theorem 2, we propose a general method for learning non-linear representations that satisfy the SO(3)-equivariance property outlined in Eq. 4. The core of our method can be abstractly referred to as TraceGrad. At the label level, it incorporates the SO(3)-invariant trace quantity $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ introduced in Theorem 1 as a supervisory signal for learning $z$, i.e., the SO(3)-invariant internal representations of $g_{\text{nonlin}}(\cdot)$ in Theorem 2. Given the attribute that invariance is preserved under non-linear operations, $z$ is encoded by $s_{nonlin}(\cdot)$ to obtain non-linear expressive capabilities. At the representation level, by taking the gradient of $z$ with respect to $\mathbf{f}$, the non-linear expressiveness of $z$ is transferred to the equivariant feature $\mathbf{v}$, while maintaining strict SO(3)-equivariance as we prove. Subsequently, merging $\mathbf{v}$ and $\mathbf{f}$, and applying $g_{\text{nonlin}}(\cdot)$ in a stacked manner can yield rich SO(3)-equivariant non-linear representations for inferring the SO(3)-equivariant regression target.
5.1 Encoding and Decoding Framework
Our encoding framework corresponds to a total of $K$ encoding modules, which are sequentially connected to form a deep encoding framework. For the $k$-th module ($1\leq k\leq K$), it first introduces an SO(3)-equivariant backbone encoder, e.g., the encoders from DeepH-E3 (Gong et al., 2023) or QHNet (Yu et al., 2023b) which is composed of recently developed equivariant operators (Thomas et al., 2018; Geiger & Smidt, 2022) like linear scaling, direct products, direct sums, Clebsch-Gordan decomposition, gated activation, equivariant normalization, and etc., to encode the physical system’s initial representations, such as spherical harmonics (Schrodinger, 1926), or representations passed from the previous neural layers, as equivariant feature $\textbf{f}^{(k)}$ in the current hidden layer. Next, we construct the feature $\textbf{v}^{(k)}=g_{\text{nonlin}}(\textbf{f}^{(k)})$, to achieve sufficient non-linear expressiveness while maintaining SO(3)-equivariance, where $g_{\text{nonlin}}(\cdot)$, $s_{nonlin}(\cdot)$ are the non-linear functions defined in Eq. 5. The function $s_{nonlin}(\cdot)$ can be implemented as any differentiable non-linear neural network module, such as feed-forward layers with non-linear activation functions like $SiLU$ and normalization operations like $Layernorm$.
For an expressive representation of a complex physical system, $\textbf{f}^{(k)}$ is usually not a feature of single degree but a direct-sum concatenation of series of components $\{\textbf{f}^{(k)l_{1}},\textbf{f}^{(k)l_{2}},\textbf{f}^{(k)l_{3}},...\}$ at multiple degrees, i.e., $L^{(k)}=\{l_{1},l_{2},l_{3},...\}$, where some components share the same degree while others differ. In this case, it becomes necessary to extend the decomposition operator, i.e., $\text{CGDecomp}(\cdot)$ in Eq. 5, to accommodate various components of degrees. Moreover, in the context of neural networks, introducing learnable parameters $W$ and more feature channels $C$ may improve the model capacity. Based on these considerations, when constructing the encoding module, the $\text{CGDecomp}(\cdot)$ operation in Eq. 5 can be expanded as $\text{CGDecomp}_{ext}(\cdot)$, and its outputs can be expanded as $\textbf{u}^{(k)}$, as shown in Eq. 6:
$\textbf{u}^{(k)}$ $=[u_{1}^{(k)},...,u_{c}^{(k)},...,u_{C}^{(k)}],$ (6)
$u_{c}^{(k)}$ $=\text{CGDecomp}_{\text{ext}}(\textbf{f}^{(k)}\otimes\textbf{f}^{(k)},0,W)=\sum_{l_{i},l_{j}\in L^{(k)},l_{i}=l_{j}}W^{c}_{ij}\cdot\text{CGDecomp}(\textbf{f}^{(k)l_{i}}\otimes\textbf{f}^{(k)l_{j}},0)$
where $W^{c}_{ij}$ represents learnable parameters, $u_{c}^{(k)}$ is the $c$ th channel of $\mathbf{u}^{(k)}$. $\text{CGDecomp}_{\text{ext}}(\cdot)$ adds learnable parameters to $\text{CGDecomp}(\cdot)$ and expands its output from a single channel to multiple channels. To further enhance the model capacity, we also expand the output of $s_{\text{nonlin}}(\cdot)$ to multiple channels as follows: $\mathbf{z}^{(k)}=[z_{1}^{(k)},\ldots,z_{c}^{(k)},\ldots,z_{C}^{(k)}]$. We define $\mathbf{v}_{c}^{(k)}=\frac{\partial z_{c}^{(k)}}{\partial\mathbf{f}^{(k)}}$, and construct the new features $\mathbf{v}^{(k)}$ by $\mathbf{v}^{(k)}=\sum_{c}\mathbf{v}_{c}^{(k)}$. It is evident that these extensions maintain the SO(3)-invariance of $\mathbf{u}^{(k)}$ and $\mathbf{z}^{(k)}$, as well as the SO(3)-equivariance of $\mathbf{v}^{(k)}$. $\mathbf{v}^{(k)}$ is combined with $\mathbf{f}^{(k)}$ in a residual manner like $\textbf{o}^{(k)}=\textbf{f}^{(k)}+\textbf{v}^{(k)}$ as the output of the $k$ th encoding module.
[…]
5.2 Training
The training loss function is shown as the following:
$\begin{split}\min_{\theta}loss&=loss_{Q}+\mu(loss_{Q},loss_{T})\cdot loss_{T},\\ loss_{Q}&=Error(\widehat{\mathbf{Q}},\mathbf{Q}^{*}),\,\,\,\,loss_{T}=Error(\widehat{\mathbf{T}},\mathbf{T}^{*})\end{split}$ (7)
where $\theta$ denotes all of the learnable parameters of our framework, $\widehat{\mathbf{Q}}$, $\widehat{\mathbf{T}}$ and $\mathbf{Q}^{*}$, $\mathbf{T}^{*}$ respectively denote the predictions and labels of $\mathbf{Q}$ and $\mathbf{T}$. In order to prevent the numerical disparity between the two loss terms and stabilize the training for both of the SO(3)-equivariant and SO(3)-invariant branches, we apply $\mu(loss_{Q},loss_{T})$, i.e., a coefficient to regularize the relative scale of the two loss terms:
$\mu(loss_{Q},loss_{T})=\lambda\cdot No\_Grad(\frac{loss_{Q}}{loss_{T}})$ (8)
where $No\_Grad(\cdot)$ denotes gradient discarding when calculating such coefficient, as this coefficient is only used for adjusting the relative scale between the two loss terms and should not itself be a source of training gradients, otherwise it would counteract the gradients from $loss_{T}$ in Eq. 7. All of the encoding and decoding modules are trained jointly by Eq. 7.
For the implementation details of our method, including the specific design of the network modules, parameter settings, and training specifics, please refer to Appendix D.
6 Experiments
6.1 Experimental Conditions
Table 1: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$ and $MAE^{H}_{cha\_b}$ metrics (meV) on the Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$) and Bilayer Bi2Se3 ($BS$) databases, where the superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets.
Methods $MG$ $MM$
$MAE$ ()($\downarrow$)( ↓ )
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.251 0.357 0.362 0.406 0.574 1.103
DeepH-E3+TraceGrad 0.175 0.257 0.228 0.285 0.412 0.808
Methods $BG^{nt}$ $BG^{t}$
DeepH-E3 (Baseline) 0.389 0.453 0.644 0.264 0.429 0.609
DeepH-E3+TraceGrad 0.291 0.323 0.430 0.198 0.372 0.406
Methods $BB^{nt}$ $BB^{t}$
DeepH-E3 (Baseline) 0.274 0.304 1.042 0.468 0.602 2.399
DeepH-E3+TraceGrad 0.226 0.256 0.740 0.384 0.503 1.284
Methods $BT^{nt}$ $BT^{t}$
DeepH-E3 (Baseline) 0.447 0.480 1.387 0.831 0.850 4.572
DeepH-E3+TraceGrad 0.295 0.312 0.718 0.735 0.755 4.418
Methods $BS^{nt}$ $BS^{t}$
DeepH-E3 (Baseline) 0.397 0.424 0.867 0.370 0.390 0.875
DeepH-E3+TraceGrad 0.300 0.332 0.644 0.291 0.302 0.674
We apply our theory and method to the electronic-structure Hamiltonian prediction task, and collect results on eight benchmark databases, i.e., Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), Bilayer Bi2Se3 ($BS$), QH9-stable (QS), and QH9-dynamic (QD). The first six databases, consisting of periodic crystalline systems with elements like C, Mo, S, Bi, Te and Se, are from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023). The last two databases are from the QH9 benchmark series (Yu et al., 2023a), composed of molecular systems with elements like C, H, O, N and F. These databases present diverse and complex challenges to the generality of a regression model. Regarding $MG$, $MM$, and QD, as their samples are prepared from an temperature environment at three-hundred Kelvin, the thermal motions lead to complex non-rigid deformations, increasing the difficulty of Hamiltonian prediction. For $BG$, $BB$, $BT$, and $BS$, the twisted structures, with an interplay of SO(3)-equivariant effects and van der Waals (vdW) force variations bring significant generalization challenges, which are further exacerbated by the absence of any twisted samples in the training sets. Besides, $BB$, $BT$, and $BS$ exhibit strong spin-orbit coupling (SOC) effects, which further increase the complexity of Hamiltonian modeling. For the $QS$ database, the ’ood’ strategy from the official settings is used to split the training, validation, and testing sets, ensuring that the atom number of samples do not overlap across the three subsets. For the $QD$ database, the ’mol’ strategy provided by Yu et al. (2023a) is applied to split the training, validation, and testing sets, ensuring that there are no thermal motion samples from the same temporal trajectory across the three subsets. The ’mol’ and ’ood’ strategies aim to assess the regression model’s extrapolation capability with respect to the number of atoms as well as the temporal trajectories, respectively. Detailed statistic information of these databases can be found in the Appendix C. Implementation details of our method for experiments on these databases are presented in Appendix D.
Table 2: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ metrics on the QH9-stable (QS) and QH9-dynamic (QD) databases respectively using ’ood’ and ’mol’ split strategies (Yu et al., 2023a). $\downarrow$ means lower values correspond to better accuracy, while $\uparrow$ means higher values correspond to better performance. The units of MAE metrics are meV, while $Sim(\psi)$ is the cosine similarity which is dimensionless.
Methods QS
MAE ()($\downarrow$)( ↓ ) $Sim(\psi)$ ()($\uparrow$)( ↑ )
$MAE^{H}_{all}$ $MAE^{H}_{diag}$ $MAE^{H}_{non\_diag}$ $MAE^{\epsilon}$
QHNet (Baseline) 1.962 3.040 1.902 17.528 0.937
QHNet+TraceGrad 1.191 2.125 1.139 8.579 0.948
Methods QD
QHNet (Baseline) 4.733 11.347 4.182 264.483 0.792
QHNet+TraceGrad 2.819 6.844 2.497 63.375 0.927
We use a comprehensive set of metrics to deeply evaluate the performance of deep learning electronic-structure Hamiltonian prediction models. The metrics can be broadly categorized into two major aspects:
• Metrics on accuracy performance. On the databases from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023), we follow Yin et al. (2024) to adopt a set of Mean Absolute Error (MAE) metrics between predicted and ground truth Hamiltonians, including $MAE^{H}_{all}$ for measuring average MAE of all samples and matrix elements, $MAE^{H}_{cha\_s}$ for measuring the MAE of challenging samples where the baseline model performs the worst, $MAE^{H}_{block}$ for measuring the MAE of different basic blocks in the Hamiltonian matrix, and $MAE^{H}_{cha\_b}$ for measuring the MAE on the most challenging Hamiltonian block where the baseline model shows the poorest performance (with the largest $MAE^{H}_{block}$). These metrics comprehensively reflect the accuracy performance, covering not only the average accuracy but also the accuracy on difficult samples and challenging blocks of the Hamiltonian matrices. On the two databases from the QH9 benchmark series, we adopt the metrics introduced by their original paper (Yu et al., 2023a), including MAE of Hamiltonian matrices, which are further subdivided into $MAE^{H}_{all}$ for measuring average MAE, $MAE^{H}_{diag}$ for measuring MAE of Hamiltonian matrix formed by an atom with itself, and $MAE^{H}_{non\_diag}$ for measuring MAE of Hamiltonian matrices formed by different atoms; as well as the MAE ($MAE^{\epsilon}$) of occupied orbital energies $\epsilon$ induced by the predicted Hamiltonians and compared to the ground truth ones, and the cosine similarity ($Sim(\psi)$) between the electronic wavefunctions $\psi$ induced by the predicted and ground truth Hamiltonians. $\epsilon$ and $\psi$ are crucial downstream physical quantities for determining multiple properties of the atomic systems as well as their dynamics, highly reflecting the application values of the Hamiltonian regression model.
• Metrics on acceleration ratios for the convergence of traditional DFT algorithms. Despite the increasing ability of deep learning models to independently handle more electronic-structure computation tasks, there are still applications with extremely high numerical precision requirements and very low tolerance for error, where traditional DFT algorithms must perform the final calculations. In such cases, the predictions from deep models can be used as initial matrices to accelerate the convergence of traditional DFT algorithms. We evaluate the acceleration ratios brought by the proposed method for the convergence of classical DFT algorithms implemented by PySCF (Sun et al., 2018). Specifically, we adopt the two metrics defined in Yu et al. (2023a): the achieved ratio calculates the number of DFT optimization steps taken when initializing with the Hamiltonian matrices predicted by the deep model compared to using initial guess methods like minao and 1e. The error-level ratio measures the number of DFT optimization steps required, starting from random initialization, to reach the same error level as the deep model’s predictions, relative to the total number of steps in the DFT process.
Table 3: The acceleration ratios of QHNet and QHNet+TraceGrad for DFT calculation. Both models are evaluated on a set of $50$ molecules chosen by Yu et al. (2023a), with the mean and standard deviation of the metrics across these samples reported. $\downarrow$ means lower values correspond to better accuracy, while $\uparrow$ means higher values correspond to better performance.
Methods Training DFT Metric Ratio
databases initialization
QHNet (Baseline) QS 1e Achieved ratio $\downarrow$ $0.400\pm 0.030$
Error-level ratio $\uparrow$ $0.620\pm 0.037$
minao Achieved ratio $\downarrow$ $0.715\pm 0.033$
Error-level ratio $\uparrow$ $0.406\pm 0.021$
QD 1e Achieved ratio $\downarrow$ $0.512\pm 0.138$
Error-level ratio $\uparrow$ $0.622\pm 0.048$
minao Achieved ratio $\downarrow$ $0.882\pm 0.217$
Error-level ratio $\uparrow$ $0.406\pm 0.066$
QHNet+TraceGrad QS 1e Achieved ratio $\downarrow$ $\textbf{0.345}\pm 0.038$
Error-level ratio $\uparrow$ $\textbf{0.685}\pm 0.037$
minao Achieved ratio $\downarrow$ $\textbf{0.647}\pm 0.061$
Error-level ratio $\uparrow$ $\textbf{0.466}\pm 0.035$
QD 1e Achieved ratio $\downarrow$ $\textbf{0.440}\pm 0.101$
Error-level ratio $\uparrow$ $\textbf{0.645}\pm 0.046$
minao Achieved ratio $\downarrow$ $\textbf{0.761}\pm 0.167$
Error-level ratio $\uparrow$ $\textbf{0.435}\pm 0.052$
6.2 Results and Analysis
We compare experimental results from two setups: the first one is the experimental results of the baseline SO(3)-equivariant regression model (Gong et al., 2023; Yu et al., 2023b) for Hamiltonian prediction, and the second one is the experimental results of extending the architecture and pipeline of the baseline model through the proposed TraceGrad method, which incorporates non-linear expressiveness into the SO(3)-equivariant features of the baseline model with the gradient operations of SO(3)-invariant non-linear features learned under the supervision of the trace targets. We choose DeepH-E3 (Gong et al., 2023) as the baseline model for databases from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023); and we choose QHNet (Yu et al., 2023b) as the baseline model for databases from the QH9 benchmark series (Yu et al., 2023a). They are the respective state-of-the-art (SOTA) methods with strict SO(3)-equivariance on the corresponding databases.
We list the results of DeepH-E3 and DeepH-E3+TraceGrad in Tables 1 for databases from the DeepH benchmark series, reporting the values of $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$, and $MAE^{H}_{cha\_b}$. The results of DeepH-E3 to be compared are copied from Yin et al. (2024). The results of DeepH-E3+TraceGrad are the average from 10 independent repeated experiments. Regarding the metric of $MAE^{H}_{block}$ for every Hamiltonian block, due to its large data volume, we just present its values for all databases from the DeepH series in Appendix 4. We use the same fixed random seed as adopted by DeepH-E3 for all random processes in experiments on these six databases. As a result, the standard deviation of the Hamiltonian prediction MAE across repeated experiments does not exceed $0.007$ meV for each of the six databases and is negligible.
From results presented in Tables 1, we could find that the proposed TraceGrad method dramatically enhances the accuracy performance of the baseline method DeepH-E3, both on average and for challenging samples and blocks, both on the non-twisted samples and the twisted samples. Specifically, on the corresponding datasets, TraceGrad lowers down the $MAE^{H}_{all}$ and $MAE^{H}_{cha}$ of DeepH-E3 with relative ratios of up to $34\%$ and $35\%$, respectively. Furthermore, from the results included in Appendix 4, TraceGrad significantly improves the performance for the vast majority of basic blocks. Particularly, for the blocks where DeepH-E3 perform the worst, TraceGrad reduces the $MAE$ ($MAE^{H}_{cha\_b}$) by a maximum of $48\%$. The leading performance on the $MG$ and $MM$ databases prepared at three-hundred Kelvin temperature demonstrates the robustness of our method against thermal motion. The high accuracy on the $BB$, $BT$, and $BS$ databases, which have strong SOC effects, indicates our method’s strong capability to model such effects. The excellent performance on the $BG^{t}$, $BB^{t}$, $BT^{t}$, and $BS^{t}$ subsets showcases the method’s superior generalization to twisted structures, which are not present in the training data. The outstanding performance on such samples highlights the good potential for studying twist-related phenomena, a hot research topic that may bring new electrical and transport properties (Cao et al., 2018; Wang et al., 2024a; He et al., 2024). Additionally, the $BG^{t}$, $BB^{t}$, $BT^{t}$, and $BS^{t}$ subsets contain significantly larger unit cells compared to the training set (see Appendix C for statistics of their sizes), yet our method still excels on these subsets as measured by the multiple MAE metrics, demonstrating its good scalability on the sizes of atomic systems it handles.
In Table 2, we present the results of QHNet and QHNet+TraceGrad under the metrics of $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ for the $QS$ and $QD$ databases. The results of QHNet to be compared are taken from their original paper (Yu et al., 2023a), and for the unification of MAE units, we convert the units of MAE from $10^{-6}$ Hartree ($E_{h}$) in the original paper to meV111 $27211.41E_{h}=$ 27211.4 meV. The results of QHNet+TraceGrad are the average from 10 independent repeated experiments. To ensure reproducibility, we use the same fixed random seeds as employed in QHNet for all random processes in the experiments on the $QS$ and $QD$ databases. As a result, the standard deviation of the Hamiltonian prediction MAE across repeated experiments is no greater than $0.009$ meV for both $QS$ and $QD$ and is also negligible.
The results presented in Table 2 demonstrate that the proposed TraceGrad method significantly enhances the accuracy of the baseline QHNet model across all metrics on the $QS$ and $QD$ databases. Specifically, TraceGrad reduces $MAE^{H}_{all}$, $MAE^{H}_{non\_diag}$, $MAE^{H}_{diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ of QHNet with relative reductions of up to 40%, 39%, 40%, 76%, and 17%, respectively, on the corresponding databases. The significant accuracy improvements on the $QS$ database, partitioned using the ’ood’ split strategy (Yu et al., 2023a) without scale overlapping among the training, validation, and testing sets, once again demonstrate the method’s strong generalization capabilities across different scales of atomic systems. Meanwhile, performance on the $QD$ database under the ’mol’ strategy (Yu et al., 2023a), which partitions the training, validation, and testing sets with samples from completely different thermal motion trajectories, highlight our method’s robustness in generalizing to new thermal motion sequences. Furthermore, the substantial improvement in the prediction accuracy of $\epsilon$, i.e., occupied orbital energies crucial for determining electronic properties such as optical characteristics and conductivity in atomic systems, and $\psi$, i.e., the electronic wavefunctions essential for understanding electron distribution and interactions, underscores the potential values of our method for applications like material design, molecular pharmacology, and quantum computing.
We also compare the acceleration ratios brought by the QHNet model and the QHNet+TraceGrad model for the convergence of classical DFT algorithms implemented by PySCF (Sun et al., 2018), with the experimental results reported in Table 3, using the achieved ratio and error-level ratio metrics introduced in Section 6.1. As observed from this table, TraceGrad also brings significant improvements to the baseline model QHNet, notably reducing the achieved ratio and enhancing the error-level ratio, indicating its substantial potential for accelerating traditional DFT algorithms.
Due to the page limits, we present three groups of supplementary experimental results in the Appendices. The first group, included in Appendix H, is the results of ablation study. These fine-grained experiments assess the individual and combined contributions of the two core mechanisms of the proposed TraceGrad method on the six databases from the DeepH series. These mechanisms are the SO(3)-invariant trace supervision mechanism (Trace) at the label level and the gradient-based induction mechanism (Grad) at the representation layer. Results indicate that each mechanism can contribute individually to the performance. Moreover, their combination provides even better performance. The second group of experiments, as included in Appendix I, test the extra time cost caused by TraceGrad compare to the baseline model. As we adopt a light-weight implementation of the non-linear modules introduced in the TraceGrad method, experimental results show that the extra time burden is slight relative to the baseline models DeepH-E3 and QHNet, indicating the good efficiency of our approach. The third group of experiments corresponds to the synergy of our method with an approximately SO(3)-equivariant methodology (Yin et al., 2024). As approximate equivariance is not the main focus of this paper, we just include these experimental results in Appendix J. The experimental results in Appendix J still demonstrate a significant accuracy improvement brought by our approach, further supporting the generality of our methodology.
7 Conclusion
We propose a theoretical and methodological framework to tackle the issue of reconciling non-linear expressiveness with SO(3)-equivariance in deep learning frameworks for physical system modeling, through deeply investigating the mathematical connections between SO(3)-invariant and SO(3)-equivariant quantities, as well as their representations. We first constructs SO(3)-invariant quantities from SO(3)-equivariant regression targets, using them to train informative SO(3)-invariant non-linear representations. From these, SO(3)-equivariant features are derived with gradient operations, achieving non-linear expressiveness while maintaining strict SO(3)-equivariance. We apply our theory and method to the challenging electronic-structure Hamiltonian prediction tasks, achieving dramatic promotions in prediction accuracy across eight benchmark databases. Experimental results demonstrate that this approach not only improves the accuracy of Hamiltonian prediction but also significantly enhances the prediction for downstream physical quantities, and also markedly improves the acceleration ratios for traditional DFT algorithms.
Ethics Statement
This work develops a representation learning method that exhibits strong non-linear expressive capabilities while strictly adhering to SO(3) equivariance. This method has demonstrated superior accuracy in predicting electronic-structure Hamiltonians and related physical quantities, showcasing its potential to accelerate research in materials science and molecular pharmacology. While we recognize that our research area has not yet revealed direct negative social or ethical implications, several issues warrant our vigilance. Currently, although our method yields accurate predictions, the decision-making processes of deep learning systems often lack transparency, hindering a comprehensive understanding of the learning outcomes and limiting our ability to gain deeper insights. We believe it is important to investigate the interpretability of such models, particularly in terms of how they apply physical knowledge in a comprehensible way. Additionally, it is crucial to continually improve the correctness and fairness of deep learning models on this area. Ensuring high-quality and diverse training data, implementing sound model designs, and performing ongoing validation and refinement are necessary to guarantee model accuracy and the broad applicability of their results.
Appendices
A Background
Definition 1.
Group. A set $G$, denoted as $G=\{\dots,g,\dots\}$, equipped with a binary operation denoted as $\cdot$, is called a group if it satisfies the following four conditions:
  1. 1.
    Closure: For any two elements $f,g\in G$, $f\cdot g\in G$.
  2. 2.
    Associativity: For all elements $f,g,h\in G$, the equality $(f\cdot g)\cdot h=f\cdot(g\cdot h)$ holds.
  3. 3.
    Existence of Identity Element: There exists a unique identity element $e\in G$ such that for all $f\in G$, $e\cdot f=f\cdot e=f$.
  4. 4.
    Existence of Inverse Element: For each element $f\in G$, there exists a unique inverse element $f^{-1}\in G$ such that $f\cdot f^{-1}=f^{-1}\cdot f=e$.
Definition 2.
SO(3) Group. The special orthogonal group $\text{SO}(3)$ is the group of all $3\times 3$ orthogonal matrices with determinant 1. Formally, it can be defined as:
$\text{SO}(3)=\{\mathbf{R}\in\mathbb{R}^{3\times 3}\mid\mathbf{R}^{T}\mathbf{R}=I,\,\det(\mathbf{R})=1\}$
where $\mathbf{R}^{T}$ denotes the transpose of $\mathbf{R}$ and $I$ is the $3\times 3$ identity matrix. The elements of $\text{SO}(3)$ represent rotations in three-dimensional Euclidean space.
Definition 3.
Group Representation. A representation of a group $G$ on a tensor space $T(V)$ is a homomorphism $\rho$ from $G$ to the general linear group $GL(T(V))$, the group of all invertible linear transformations on $T(V)$. Here, $T(V)$ represents the tensor space associated with the vector space $V$, encompassing all tensors that can be formed from elements of $V$. The homomorphism $\rho$ can be formalized as:
$\rho:G\to GL(T(V))$
such that for all $g_{1},g_{2}\in G$,
$\rho(g_{1}g_{2})=\rho(g_{1})\rho(g_{2})$
and $\rho(e)=I$, where $e$ is the identity element of $G$, and $I$ is the identity transformation on $T(V)$.
Definition 4.
SO(3) Group Representation. A representation of the special orthogonal group $\text{SO}(3)$ on a tensor space $T(V)$ is a homomorphism $\rho:\text{SO}(3)\to GL(T(V))$, where $GL(T(V))$ denotes the group of all invertible linear transformations on the tensor space $T(V)$. This representation can be expressed using Wigner-D matrices $\mathbf{D}^{l}(\mathbf{R})$, which are defined for a given degree $l$ corresponding to the representation of the rotation matrix $\mathbf{R}\in\text{SO}(3)$:
$\rho(\mathbf{R})=\mathbf{D}^{l}(\mathbf{R}).$
Definition 5.
Equivariance with Respect to a Group. Let $G$ be a group, and let $\rho_{T(V)}:G\to GL(T(V))$ and $\rho_{T(W)}:G\to GL(T(W))$ be representations of $G$ on tensor spaces $T(V)$ and $T(W)$, respectively. A map $f:T(V)\to T(W)$ is said to be equivariant with respect to the group $G$ if the following condition holds:
$f(\rho_{T(V)}(g)\circ v)=\rho_{T(W)}(g)\circ f(v)\quad\text{for all }v\in T(V)\,\,and\,\,g\in G.$
where $\circ$ generally denotes the operation defined on the tensor space.
Definition 6.
Invariance with Respect to a Group. Let $G$ be a group, and let $\rho_{T(V)}:G\to GL(T(V))$ be a representation of $G$ on a tensor space $T(V)$. A function $f:T(V)\to T(W)$ is said to be invariant under the group $G$ if the following condition holds:
$f(\rho_{T(V)}(g)\circ v)=f(v)\quad\text{for all }v\in T(V)\,\,and\,\,g\in G.$
This definition indicates that the function $f$ remains unchanged under the action of the group $G$.
B Proofs of Theorems
Proof of Theorem 1.
Under an SO(3) rotation represented by the rotational matrix $\mathbf{R}$, $\mathbf{Q}=\mathbf{Q}^{l_{p}\otimes l_{q}}$ is transformed as $\mathbf{Q}(\mathbf{R})$:
$\mathbf{Q}(\mathbf{R})=\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger},$
where $\mathbf{D}^{l_{p}}(\mathbf{R})$ and $\mathbf{D}^{l_{q}}(\mathbf{R})$ are the Wigner-D matrices for the degrees of $l_{p}$ and $l_{q}$, respectively, corresponding to the rotation $\mathbf{R}$.
The conjugate transpose of the transformed quantity is:
$\mathbf{Q}(\mathbf{R})^{\dagger}=\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}.$
Using the cyclic property of the trace, which states that the trace of a product of matrices remains unchanged under cyclic permutations (i.e., $tr(ABC)=tr(BCA)=tr(CAB)$), and combining the properties that $\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}=\mathbf{I}$ and $\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger}=\mathbf{I}$, we can rearrange the terms inside the trace as follows:
$\mathbf{T}(\mathbf{R})=tr(\mathbf{Q}(\mathbf{R})\cdot\mathbf{Q}(\mathbf{R})^{\dagger})=tr((\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger})\cdot(\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}))$
$=tr(\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger})=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R}))=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\mathbf{T}.$
Therefore, $\mathbf{T}=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is invariant under SO(3) transformations, its SO(3)-invariance is proved. ∎
Proof of Theorem 2.
Under the given condition, the input feature $\mathbf{f}$ in direct-sum state is SO(3)-equivariant, meaning that under an SO(3) rotation represented by $\mathbf{R}$, it transforms as follows:
$\mathbf{f}(\mathbf{R})=\mathbf{D}^{l}(\mathbf{R})\cdot\mathbf{f}$
where $\mathbf{D}^{l}(\mathbf{R})$ is the Wigner-D matrix corresponding to degree $l$.
First, according to group theory, $u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ is an SO(3)-invariant scalar as the degree-zero component from the Clebsch-Gordan decomposition is invariant under rotations. Since applying a non-linear operation to an SO(3)-invariant quantity does not change its invariance, $z=s_{\text{nonlin}}(u)$ is also SO(3)-invariant, independent to the specific form of $s_{\text{nonlin}}(\cdot)$. It formally holds that:
$z(\mathbf{R})=z$ (9)
Next, we apply the chain rule in Jacobian form. Considering $\mathbf{f}(\mathbf{R})$ is in the form of a column vector, to facilitate the application of the chain rule in vector form, we first transpose it into a row vector $\mathbf{f}(\mathbf{R})^{T}$, then differentiate:
$\frac{\partial z(\mathbf{R})}{\partial\mathbf{f}^{T}(\mathbf{R})}=\frac{\partial z}{\partial\mathbf{f}^{T}(\mathbf{R})}=\frac{\partial z}{\partial\mathbf{f}^{T}}\frac{\partial\mathbf{f}^{T}}{\partial\mathbf{f}(\mathbf{R})^{T}}=\frac{\partial z}{\partial\mathbf{f}^{T}}\cdot\mathbf{D}^{l}(\mathbf{R})^{-1}=\frac{\partial z}{\partial\mathbf{f}^{T}}\cdot\mathbf{D}^{l}(\mathbf{R})^{T}$ (10)
Here we utilize the property that $\mathbf{D}^{l}(\mathbf{R})^{-1}=\mathbf{D}^{l}(\mathbf{R})^{T}$  [^2] In Theorem 1 and Theorem 2, the Wigner-D matrices are in the complex and real fields, respectively, since the target quantity may be complex, whereas the internal representations of neural networks are typically in the real field. Nonetheless, neural network representations in the real field can still predict complex-valued targets with SO(3)-equivariance. Previous literature (Gong et al., 2023) has provided mechanisms for converting the network outputs in the real field into regression targets with real and imaginary parts.. Since the representations of neural networks are generally real numbers, the corresponding Wigner-D matrix is also real unitary.
Finally, we transpose the result back to a column vector:
$\mathbf{v}(\mathbf{R})=g_{nonlin}(\mathbf{f}(\mathbf{R}))=(\frac{\partial z(\mathbf{R})}{\partial\mathbf{f}^{T}(\mathbf{R})})^{T}=(\frac{\partial z}{\partial\mathbf{f}^{T}}\cdot\mathbf{D}^{l}(\mathbf{R})^{T})^{T}=\mathbf{D}^{l}(\mathbf{R})\cdot\frac{\partial z}{\partial\mathbf{f}}=\mathbf{D}^{l}(\mathbf{R})\cdot\mathbf{v}$ (11)
This proves that $g_{\text{nonlin}}(\cdot)$ is an SO(3)-equivariant non-linear operator: when applying its non-linearity to a SO(3)-equivariant feature $\mathbf{f}$, the output feature $\mathbf{v}$ remains SO(3)-equivariant. ∎
C Information of Experimental Databases
Figure 2: Illustration of a Hamiltonian matrix. Each Hamiltonian matrix contains multiple basic blocks, with the $\mathbf{Q}=\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 3 corresponding to a basic block $\mathbf{H}^{l_{p}\otimes l_{q}}$ here.
In this part, we provide detailed information about the experimental databases, including the statistical information of the six databases from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023) and the two databases from the QH9 benchmark series (Yu et al., 2023a), listed in Table 4 and Table 5, respectively. Additionally, we visualize two types of challenging testing samples: samples with non-rigid deformation from thermal motions, as well as the bilayer samples with interlayer twists, which are shown in Fig. 3 and Fig. 4, respectively.
D Implementation Details
The hardware environment for our experiments is a server cluster equipped with Nvidia RTX A6000 GPUs, each with 48 GiB of memory. Other experimental details may differ across the DeepH and QH9 benchmark series, which we will describe separately.
E Implementation Details on the DeepH Benchmark Series
The software environment used is Pytorch 2.0.1 for experiments on the six crystalline databases from the DeepH benchmark series. When combining the proposed TraceGrad method with the DeepH-E3 architecture, the implementation of DeepH-E3 is based on the project  [^3] https://github.com/Xiaoxun-Gong/DeepH-E3 provided by Gong et al. (2023), keeping the architecture and hyperparameter configurations consistent with their setup. In our framework, we use the same number of encoding modules $K$ as DeepH-E3, which is set to $3$. In each encoding module, we apply the $g_{nonlin}(\cdot)$ module proposed in Section 4 and 5 to each SO(3)-equivariant edge feature, enabling non-linear expressiveness. We set the number of channels for the rotationally SO(3)-invariant feature $\textbf{u}^{(k)}$ $(1\leq k\leq 3)$ to $1024$. The neural network module $s_{nonlin}(\cdot)$ within $g_{nonlin}(\cdot)$ is implemented as a three-layers fully-connected module: the input size is set to $1024$, consistent with $\textbf{u}^{(k)}$; the hidden layer size is also $1024$, with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; and the output layer size (i.e., the dimensionality of $\textbf{z}^{(k)}$) is set to be equal to the number of basic blocks for a Hamiltonian matrix, which is $25$ for $MG$ and $BG$, $49$ for $MM$, and $196$ for $BB$, $BT$, and $BS$. It is worth noting that, while $s_{nonlin}(\cdot)$ can be implemented as any differentiable neural network module, we here implement it as a simple fully-connected module. This decision is made to avoid adding significant computational burden to the whole network. Meanwhile, as DeepH-E3 already incorporates complex graph network mechanisms for information aggregation and message-passing, there is no need for $s_{nonlin}(\cdot)$ to be overly complex. Its role is focusing on to filling in the gaps left by the existing equivariant mechanisms in DeepH-E3: to introduce a non-linear mapping mechanism that maintains equivariance, thereby activating and unleashing the expressive power of the overall network architecture through non-linearity. The SO(3)-equivariant decoder we adopt is the same as that of DeepH-E3; The SO(3)-invariant decoder we adopt is a four-layers fully-connected module: the input size is $3$ $(K)$ times of the dimensionality of $\textbf{z}^{k}$, e.g., $75$ for $MG$; the hidden layers have $1024$ neurons with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; the size of the output layer is the number of basic blocks for a Hamiltonian matrix. Since each basic block of the Hamiltonian matrix can compute a trace, the total number of trace variables corresponds to the number of basic blocks. Regarding the error metric in the loss function Eq. 7, for the first term, we follow DeepH-E3 to use MSE (Mean Squared Error); for the second term, we choose between MSE and MAE based on performance on the validation sets, ultimately selecting MAE. $\lambda$ in the training loss function is set according to parameter selection on the validation sets, searching from $\{0.1,0.2,...,1.0\}$. Here, we aim to obtain a more general parameter setting for $\lambda$ on crystalline structures, and thus we determined $\lambda$ based on the overall performance on the validation sets of the six crystalline databases and the searched value is $0.3$. Other hyper-parameters and configurations are the same as DeepH-E3 (Gong et al., 2023): the initial learning rates for experiments on the $MG$, $MM$, $BG$, $BB$, $BT$, and $BS$ databases are set to $0.003$, $0.005$, $0.003$, $0.005$, $0.004$, and $0.005$, respectively; the training batch size is set as $1$; the optimizer is chosen as Adam; the scheduler is configured as a slippery slope scheduler.
F Implementation Details on the QH9 Benchmark Series
The software environment used is Pytorch 1.11.0 for experiments on the two molecular databases from the QH9 benchmark series. When combining the proposed TraceGrad method with the QHNet architecture, the implementation of QHNet is based on the project  [^4] https://github.com/divelab/AIRS provided by Yu et al. (2023b), keeping the architecture and hyperparameter configurations consistent with their setup. In our framework, we use the same number of encoding modules as QHNet: $5$ node feature encoding modules and $2$ edge feature encoding modules. We opt to apply the $g_{nonlin}(\cdot)$ module proposed in Section 4 and 5 to each SO(3)-equivariant edge feature. We set the number of channels for $\textbf{u}^{(k)}$ $(1\leq k\leq 2)$ as $1024$. The neural network module $s_{nonlin}(\cdot)$ within $g_{nonlin}(\cdot)$ is implemented as a three-layers fully-connected module: the input size is set to $1024$, consistent with $\textbf{u}^{(k)}$, the hidden layer size is also $1024$, with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism, and the output layer size (i.e., the dimensionality of $\textbf{z}^{(k)}$) is set to be equal to the number of basic blocks for a Hamiltonian matrix, which is $36$ for $QS$ and $QD$ databases. The SO(3)-equivariant decoder we adopt is the same as that of QHNet; the SO(3)-invariant decoder we adopt is a four-layers fully-connected module: the input size is $2$ $(K)$ times of the dimensionality of $\textbf{z}^{(k)}$, e.g., $72$ for $QS$ and $QD$; the hidden layers have $1024$ neurons with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; the size of the output layer is the number of basic blocks for a Hamiltonian matrix. Regarding the error metric in the loss function Eq. 7, for the first term, we follow QHNet to use a combination of MSE and MAE; for the second term, we choose between MSE and MAE based on performance on the validation sets, ultimately selecting MAE. $\lambda$ in the training loss function is set according to parameter selection on the validation sets, searching from $\{0.1,0.2,...,1.0\}$. Here, we aim to obtain a more general parameter setting for $\lambda$ on molecular structures, and thus we determined $\lambda$ based on the overall performance on the validation sets of the two molecular databases and the searched value is $0.2$. Other hyper-parameters and configurations are the same as QHNet: the initial learning rates for all experiments are set as $5\times 10^{-4}$, the training batch size is set as $32$, the optimizer is selected as AdamW, and a learning rate scheduler is implemented. To be more specific, the scheduler gradually increases the learning rate from 00 to a maximum value of $5\times 10^{-4}$ over the first $1,000$ warm-up steps. Subsequently, the scheduler linearly reduces the learning rate, ensuring it reaches $1\times 10^{-7}$ by the final step.
G Visualization of Block-level MAE Statistics
As shown in Fig. 2, each Hamiltonian matrix consists of numerous basic block, with each basic block representing the direct product of two degrees. Here, we follow Yin et al. (2024) to measure the MAE performance of deep models on each basic block, denoted as $MAE^{H}_{block}$. The values of $MAE^{H}_{block}$ for the two setups, i.e., DeepH-E3 and DeepH-E3+TraceGrad, on different blocks of the Hamiltonian matrix for six databases from the DeepH benchmark series are illustrated in Fig. 5 and 6. Fig. 5 presents the results for monolayer structures, while Fig. 6 focuses on bilayer structures. From these figures, it can be observed that our method, TraceGrad, brings significant accuracy improvements over the baseline method, DeepH-E3, across the vast majority of blocks of the Hamiltonian matrices, particularly on blocks where DeepH-E3 struggles with lower accuracy.
Table 4: Statistical information of the six benchmark databases, i.e., Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), Bilayer Bi2Se3 ($BS$), from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023). SOC: effects of Spin-Orbit Coupling. $m$: number of samples in the current dataset; $a_{max}$: maximum number of atoms from a unit cell in the current dataset. $a_{min}$: minimum number of atoms from a unit cell in the current dataset. $nt$: non-twisted samples. $t$: twisted samples.
Statistic Types MG MM BG BB BT BS
Elements C Mo, S C Bi Bi, Te Bi, Se
SOC weak weak weak strong strong strong
Training ($nt$) $m$ 270 300 180 231 204 231
$a_{max}$ 72 75 64 36 90 90
$a_{min}$ 72 75 64 36 90 90
Validation ($nt$) $m$ 90 100 60 113 38 113
$a_{max}$ 72 75 64 36 90 90
$a_{min}$ 72 75 64 36 90 90
Testing ($nt$) $m$ 90 100 60 113 12 113
$a_{max}$ 72 75 64 36 90 90
$a_{min}$ 72 75 64 36 90 90
Testing ($t$) $m$ 9 4 2 2
$a_{max}$ 1084 244 130 190
$a_{min}$ 28 28 70 70
Table 5: Statistical information of the two benchmark databases, QH9-stable ($QS$) and QH9-dynamic ($QD$), from the QH9 benchmark series (Yu et al., 2023a). The $QS$ database is split using the ’ood’ strategy, while the $QD$ database is split using the ’mol’ strategy. $m$: number of samples in the current dataset. $a_{max}$: maximum number of atoms for a sample in the current dataset. $a_{min}$: minimum number of atoms for a sample in the current dataset.
Statistic Types QS QD
Elements C, H, O, N, F C, H, O, N, F
Training $m$ 104,001 79,900
$a_{max}$ 20 19
$a_{min}$ 3 10
Validation $m$ 17,495 9,900
$a_{max}$ 22 19
$a_{min}$ 21 10
Testing $m$ 9,335 10,100
$a_{max}$ 29 19
$a_{min}$ 23 10
Figure 3: Visualization of testing samples exhibiting non-rigid deformations due to thermal motions
Figure 4: Visualization of testing samples with interlayer twists.
Figure 5: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the Monolayer Graphene ($MG$) and Monolayer MoS2 ($MM$) databases.
Figure 6: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the non-twisted (marked with superscripts $nt$) and twisted (marked with superscripts $t$) testing subsets of Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), and Bilayer Bi2Se3 ($BS$).
H Ablation Study
We conduct fine-grained ablation study on the six databases from DeepH benchmark series, comparing results from the four setups:
• DeepH-E3 (Gong et al., 2023): the baseline model.
• DeepH-E3+Trace: this experimental setup, an ablation term, only implements half part of our method. Specifically, it extends the architecture of DeepH-E3 by adding our SO(3)-invariant encoding and decoding branches and using the trace quantity $\mathbf{T}=\text{tr}(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\text{tr}(\mathbf{H}^{l_{p}\otimes l_{q}}\cdot(\mathbf{H}^{l_{p}\otimes l_{q}})^{\dagger})$ to train them. As for ablation study, this setup does not include the gradient-based mechanism delivering non-linear expressiveness from SO(3)-invariant features to encode SO(3)-equivariant features; instead, it directly uses the SO(3)-equivariant features outputted by DeepH-E3 for Hamiltonian regression. In this configuration, the SO(3)-invariant branches only contribute indirectly during the training phase by backpropagating the supervision signals from the trace quantity to the earlier layers.
• DeepH-E3+Grad: this setup is also an ablation term and implements the other half part of our method in contrast to the previous ablation term. Specifically, it incorporates our SO(3)-invariant encoder branch as well as the gradient-induced operator to deliver non-linear expressiveness from SO(3)-invariant features to encode SO(3)-equivariant features. As for ablation study, this setup continues to use the single-task training pipeline of DeepH-E3, supervised only with the Hamiltonian label without joint supervised training through the trace of Hamiltonian.
• DeepH-E3+TraceGrad: this is a complete implementation of our framework extending beyond of the architecture and training pipeline of DeepH-E3, at the label level, we introduce the trace quantity to guide the learning of SO(3)-invariant features; Meanwhile, at the representation level, we leverage the gradient operator to yield SO(3)-equivariant non-linear features for Hamiltonian prediction.
Table 6: Ablation study MAE results (meV) on the Monolayer Graphene ($MG$) and Monolayer MoS2 ($MM$) databases. $\downarrow$ means lower values of the metrics correspond to better accuracy.
Methods $MG$ $MM$
$MAE$ ()($\downarrow$)( ↓ )
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.251 0.357 0.362 0.406 0.574 1.103
DeepH-E3+Trace 0.230 0.344 0.348 0.378 0.537 1.091
DeepH-E3+Grad 0.185 0.269 0.258 0.308 0.453 0.924
DeepH-E3+TraceGrad 0.175 0.257 0.228 0.285 0.412 0.808
Table 7: Ablation study MAE results (meV) on the Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), and Bilayer Bi2Se3 ($BS$) databases. The superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets. $\downarrow$ means lower values of the metrics correspond to better accuracy.
Methods $BG^{nt}$ $BG^{t}$
$MAE$ ()($\downarrow$)( ↓ )
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.389 0.453 0.644 0.264 0.429 0.609
DeepH-E3+Trace 0.362 0.417 0.593 0.251 0.401 0.480
DeepH-E3+Grad 0.320 0.356 0.511 0.222 0.389 0.446
DeepH-E3+TraceGrad 0.291 0.323 0.430 0.198 0.372 0.406
Methods $BB^{nt}$ $BB^{t}$
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.274 0.304 1.042 0.468 0.602 2.399
DeepH-E3+Trace 0.259 0.285 0.928 0.429 0.570 1.782
DeepH-E3+Grad 0.243 0.272 0.824 0.406 0.542 1.431
DeepH-E3+TraceGrad 0.226 0.256 0.740 0.384 0.503 1.284
Methods $BT^{nt}$ $BT^{t}$
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.447 0.480 1.387 0.831 0.850 4.572
DeepH-E3+Trace 0.406 0.462 1.239 0.784 0.812 4.520
DeepH-E3+Grad 0.342 0.365 0.750 0.742 0.786 4.463
DeepH-E3+TraceGrad 0.295 0.312 0.718 0.735 0.755 4.418
Methods $BS^{nt}$ $BS^{t}$
$MAE_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.397 0.424 0.867 0.370 0.390 0.875
DeepH-E3+Trace 0.382 0.397 0.843 0.351 0.367 0.838
DeepH-E3+Grad 0.343 0.365 0.696 0.324 0.339 0.746
DeepH-E3+TraceGrad 0.300 0.332 0.644 0.291 0.302 0.674
Experimental results of the four setups are listed in Table 6 and 7. Table 6 presents the results for monolayer structures, while Table 7 focuses on bilayer structures. From the results of ablation terms, we can obtain a more fine-grained experimental analysis. By comparing among the results of DeepH-E3, DeepH-E3+Trace, DeepH-E3+Grad, and DeepH-E3+TraceGrad, we can conclude that the two core mechanisms of our method, i.e., the SO(3)-invariant trace supervision mechanism (Trace) at the label level as well as the gradient-based induction mechanism (Grad) at the representation layer, can contribute to the performance individually. Moreover, their combination provides even better performance. This is because, on one hand, with the gradient-based induction mechanism as a bridge, the non-linear expressiveness of SO(3)-invariant features learned from the trace label can be transformed into the SO(3)-equivariant representations during inference; on the other hand, with trace label, the SO(3)-invariant network branch has a strong supervisory signal, enabling it to learn the intrinsic symmetry and complexity of the regression targets, enhancing the quality of SO(3)-invariant features and ultimately benefits the encoding of SO(3)-equivariant features. The value of such complementarity has been fully demonstrated in the experimental results.
I Discussion on Time Costs
When the atomic system is small and any two atoms are within the cutoff radius of each other such that the Hamiltonian is defined, the number of Hamiltonian matrices is approximately proportional to $N^{2}$, where $N$ is the number of atoms in the system. Yet, as the atomic system grows, the system exhibits a dual nature of locality and non-locality. Locality refers to the fact that, when $N$ becomes large, the number of atoms within the cutoff radius of each atom remains roughly constant, and thus the growth in the number of Hamiltonian matrices asymptotically scales as $O(N)$. On the other hand, the non-locality arises from the fact that, although Hamiltonian matrices are defined only for pairs of atoms within each other’s cutoff radius, atoms outside the cutoff radius may still exert long-range effects on the Hamiltonian matrices within the cutoff radius
The baseline models we select, whether DeepH-E3 (Gong et al., 2023) or QHNet (Yu et al., 2023b), are SO(3)-equivariant graph neural network models with efficient information aggregation and message-passing mechanisms, which cleverly balance the locality and non-locality of the system. As a result, the computational complexity asymptotically scales as $O(N)$ as $N$ grows to large. This is a significant efficiency advantage compared to the $O(N^{3})$ computational complexity of traditional DFT methods, enabling efficient and low-cost prediction of electronic structures for large-scale atomic systems. Our method improves the Hamiltonian prediction accuracy of DeepH-E3 and QHNet while maintaining this advantage of asymptotic complexity. The neural module $g_{nonlin}(\cdot)$ we introduce efficiently updates each edge feature $\textbf{f}^{(k)}$ $(1\leq k\leq K)$ from the baseline models to obtain the corresponding non-linear SO(3)-equivariant feature $\textbf{v}^{(k)}$, ensuring that the computational complexity remains asymptotically $O(N)$, with good potential to handle large atomic systems at acceptable computational costs.
We test the inference time costs of the DeepH-E3 and DeepH-E3+TraceGrad models on an Nvidia RTX A6000 GPU, using the testing sets from two databases, namely $MG$ and $MM$, as examples; and test the inference time costs of the QHNet and QHNet+TraceGrad models on $QS$ and $QD$, on the same device. The batch size is set to 1 for all tests on inference time. Each sample is inferred 50 times, and the average time per inference is recorded. When using DeepH-E3, the average inference time per sample (s/sample) for $MG$ and $MM$ are as follows: $0.247$ and $0.256$, respectively; with DeepH-E3+TraceGrad, the average inference times are $0.264$ and $0.274$, correspondingly. When using QHNet, the average inference time per sample (s/sample) for $QS$ and $QD$ are as follows: $0.233$ and $0.174$, respectively; with QHNet+TraceGrad, the average inference times are $0.248$ and $0.187$, correspondingly. As we can see, the increases in computational costs due to the TraceGrad module are slight relative to that of the baseline models. Considering the significant accuracy improvements brought by the TraceGrad method as reported in Section 6.2, we believe that this minor increase in computational cost is acceptable and worthwhile in practical applications.
J Empirical Study on Combining our Method with Approximately SO(3)-equivariance Framework
While non-strict SO(3)-equivariance, which may limit the depth of theoretical exploration, is not the main focus of this study aiming at bridging rigorous SO(3)-equivariance with the non-linear expressive capabilities of neural networks, considering that they remain of interest in a few numerical computation applications where precision is highlighted over strict equivariance, we also conduct empirical study combining our method with approximately SO(3)-equivariant techniques. Taking the Monolayer MoS2 ($MM$) database as a case study, we evaluate the performance of combining our trace supervision and gradient induction method (TraceGrad) with the an approximately equivariant approach HarmoSE (Yin et al., 2024). We here take HarmoSE as the backbone encoder, and yields features by TraceGrad to enrich its representations. The experimental results in Table 8 and Fig. 7 demonstrate that TraceGrad significantly enhances the accuracy of HarmoSE, surpassing DeepH-2 and achieving SOTA results. Both DeepH-2 and HarmoSE sacrificed strict SO(3)-equivariance to fully release the expressive capabilities of graph Transformers, aiming for the ultimate in prediction accuracy. Despite this, our method still manages to significantly exceed their accuracy, further confirming the superiority and generality of our method in learning expressive representations of physical systems.
Figure 7: Comparison on the $MAE^{H}_{block}$ metric for HarmoSE and HarmoSE+TraceGrad on the $MM$ database.
Table 8: MAE results (meV) for DeepH-2, HarmoSE, and HarmoSE+TraceGrad on the $MM$ database. $\downarrow$ means lower values of the metrics correspond to better accuracy. The results of the compared methods are taken from the corresponding literature (Wang et al., 2024b; Yin et al., 2024), where the empty items are due to the data not being provided in the original paper.
Methods $MM$
$MAE$ ()($\downarrow$)( ↓ )
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-2 (Wang et al., 2024b) 0.21 - -
HarmoSE (Yin et al., 2024) 0.233 0.293 0.406
HarmoSE+TraceGrad 0.178 0.228 0.296
K Future Work
In future research, various extensions are conceivable across theoretical, methodological, and application fields:
First, from a theoretical and methodological perspective, we can generalize SO(3) group symmetry to a broader range and more complex groups in deep learning research. This will enable deep learning frameworks to accommodate richer mathematical structures and physical quantities. Second, from the application perspective, the current methodology extends beyond predicting electronic-structure Hamiltonians and is capable of generally predicting multiple physical quantities and properties that are equivariant. Furthermore, as our approach offers a promising general solution to the critical dilemma between equivariance and non-linear expressiveness for the deep learning paradigm, the methodology is not confined to physical research alone and may find applications in areas such as robotics, autonomous vehicles, and motion tracking systems. For instance, in the vision tasks of autonomous vehicles (Wang et al., 2023), where there is a high demand for robustness against coordinate transformations of cameras, the mainstream approach has been to achieve approximate SO(3)-equivariance through data augmentation. However, this does not guarantee absolute reliability and safety. Our work holds great potential to construct deep models with high generalization performance under the premise of strict SO(3)-equivariance, potentially revolutionizing techniques in such tasks. Looking forward to future endeavors, the foundational and pivotal role of this work is clear.